A Bibliography of Publications in Software Process: Improvement and Practice

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 January 2018 Version 1.00

Title word cross-reference

2 [BM06]. ^{SM} [KCF^{+}96].

0 [Ben05]. 0-87389-594-0 [Ben05].

15504 [Col07, Jun03, Rou03, RT07, SER^{+}97, Sta00].

2 [JHGEE01]. 2003 [WS04]. 2004 [Din06, PR05, vOW05]. 2005 [BM06, MP06, Rou07]. 2006 [Wer07]. 2007 [AMK08]. 29th [Pau96].

3rd [vOW05].

4 [AGV04]. 4everedit [MRS06]. 4th [WS04].

5th [PR05].

6th [MP06].

9 [Ano05]. 9000 [SMH96]. 9001 [Oul96]. 93 [Ano05].

Ågerfalk [Ano09]. Accelerating [IPB01]. accelerator [IPB01]. accessibility [Ako09]. Accurate [MWE^{+}09]. achieve [DC08, MŠWM06]. achieved [SB06]. achievement [Wal03]. Achieving [Ben05]. acquisition [DDT01, Häb04, MY06]. across [JS05]. active [Pen00]. activities [LT98, LMSW08]. adaptable [FR05]. adaptation [Byg05]. Addressing [LT98, PhD09a, Dal04b]. adept [CRa08]. Adopting [LVL^{+}06, BDIM06]. Adoption [BSN07, Bos05b, TGC^{+}08]. Advanced [BT98]. Advances [CGA07]. Aerospace [LT98]. Agent [SCFR06, XOZ^{+}07]. Agent-based [SCFR06]. agile [KL08, SA07, SS07, TGC^{+}08]. Agility [Dal07a, Ket07, ABM07, KL08]. agreement [EBS96]. AHP [Jun01]. Airways [ABEL97]. Aligning [ABS05]. alignment [Tax05]. alive [DR03b]. all-pervasive [Dal04b]. allocation [oSWR07]. alone [Dal05a]. AMETIST [Tho98]. among [SMH96]. analyse [BHO07]. Analysing [WGHS01]. Analysis [DP08, ASKM09, AEP08, BH02, Car07, CVO^{+}04, CWZ06, LVL^{+}06, LM04, LR97, OBM05, RFoSS08, RH01, RA05, Sca00, Sof05, WMR04, APK06]. Analyzing [PAER07]. application [CR04, GKP09, NKM08, WGHS01]. applications [NE05, vdWBSV06]. Applied [Rom96, RCHSD00]. Applying [NNR96, XOZ^{+}07]. appraisal [WCM^{+}07]. Appreciative [HNOS09]. Approach [Lar08, Li07, AGJ^{+}00, CR09, CBK06, CZS06, FR05, GGdASA06, Har05, LVL^{+}06, Moi98, Tax05, TD08]. Approaches [TM97, Tom96, Dal08]. architectural [AHI^{+}05]. Architecture [Ban96, DP96, ORM03, FvGB03, GO05, HSBR03, LMSW08, NS07, SWG^{+}07]. area [DR05]. areas [WMM05]. articles [Din06]. artifacts [LGA07]. Aspects [FD98, Pot09]. ASQ [Ben05]. Assessing [MBL07, Col07, Pad02]. Assessment [Li07, ME07, Miy96, AGJ^{+}00, CRa08, CFS01, CGA07, DD03, DHP^{+}07, HRW97, KSK00, MV07, Oja08, Oja09, Rou07, TGC^{+}08]. Assessment-based [ME07]. Assessments [SS96, Jun01, MC09, MRH^{+}09, SER^{+}97, vWVS06]. Assessor [EBS96]. Associates [Miy96]. assurance [BRMG09, LK97, LGA07, RMCGSR07]. Attaining [Dal09a]. attribute [Jun01].

August [vOW05]. Automated [Ram06]. Automotive [CRa08, MRH^{+}09, FFL^{+}07]. aviation [LVL^{+}06].

Barriers [DP96]. base [SS01]. based [AHI^{+}05, Ano05, BDVD05, BJ07, BDR96, Can04, CC96, CWZ06, CHE05a, FFL^{+}07, GKP09, HW04, IPB01, Jun01, KH08, LSK^{+}06, MRS06, ME07, MERO08, Moi98, Pen00, RRT04, RM00, SCFR06, Sta00, TM01, WGJ^{+}08, YB07, vdWBSV06, vWVS06]. Basel [DHP^{+}07]. be [JPH07]. benchmarking [DN07]. Benefits [Ano09, ÓOÅF09, HW04]. best [ME07, Sat97]. between [FvGB03, KB03, MLC^{+}08, RSB05, SM03]. BiDefect [GWY^{+}09]. biggest [Šmi06]. BL [FSC^{+}08]. black [ANB07]. black-box [ANB07]. Book [Ben05]. Boston [vOW05]. both [MR00]. bottom [Bro04]. bottom-up [Bro04]. box [ANB07]. BPEL [Car07]. Brazil [MRW09]. Brian [Ano09]. bridge [SM03]. bridging [KB03]. British [ABEL97]. building [FR05, JSZ97]. Business [GRSZ07, LSK^{+}06, ABS05, BWRC05, BDIM06, DN07, DKC00, MY06, RSB05, RBW07, Sof05]. Business-oriented [LSK^{+}06].

calibration [GKP09]. can [Dal06a].

Capability

[Dal04a, KCF^{+}96, RM00, Vak97, Col07, Dal04b, Gar97, JH03, Jun03, Sai03, Wal03]. Capturing [BWRC05, FvGB03]. cardinality [CHE05a]. cardinality-based [CHE05a]. Case [BCQ02, GGdASA06, MT00, ANB07, ASKM09, AR07, BHJ06, BD00, CLRW00, Ger03, HSBR03, JC09, MŠ08, MR06, NKM08, NNR96, Oja08, PDL08, Vak97, WMR04, Wal03]. Catalogues [CFQ08]. CBSE [TGG08]. Celebrating [Dal07b]. center [MLN^{+}03]. centered [GGB03]. Centre [ESM96]. centred [JSZ97]. Certification [LGA07]. challenge [Kil97]. challenges [CM06, Dam03, TBD08]. Change [Ano98, Dal07c, KM05, Kil97, Lin98, Moi98, PhD09b, Qin07, SM98]. change-prediction [Lin98]. changes [Sof05, WMR04]. changing [HBK06]. characteristics [TT06]. characterization [Can04]. Characterizing [RSMM05, TNT^{+}05]. Checks [JC09]. Chinese [CLM^{+}08, MLC^{+}08]. civil [LVL^{+}06]. classifying [GO05]. CMM [RM00, AGV04, BD00, Can04, HW04, IPB01, LSK^{+}06, Oul96]. CMM-based [HW04, IPB01]. CMMI [BDVD05, CFQ08, RT07, WMM05]. co [Dal06a]. co-exist [Dal06a]. coherence [CFS01]. Collaboration [PL03, Pen00, SB06]. collaborative [FGGM04, dAB07]. Combining [KL08, WH02]. commerce [DC08]. commitment [dAB07]. Common [Häb04]. Commonality [RA05, OBM05]. communication [MLC^{+}08]. Communications [Hum96, PS96b, PS96c]. companies [FG06, GGdASA06, HRG00, JC09, Šmi06, TGC^{+}08, Tho98, vWVS06, Ben05]. company [AEH^{+}08, BHJ06, MT97, MD06, Ric01]. competence [RTM09]. complex [Dal06e]. Complexity [Car07, Dal04b, HRG00]. compliance [Ben05, BDIM06]. component [KH08, Ram06, Sch07]. component-based [KH08]. components [CLM^{+}08, ŞCVB05, TNT^{+}05]. composable [ŞCVB05]. composition [GL08]. comprehensive [AGJ^{+}00]. computer [KB03]. concept [DLW06]. concepts [SS01]. Conceptual [Lep06, EC03]. Conchúir [Ano09]. concordance [DHP^{+}07]. conducting [DDA05]. Conference [Pau96, vOW05]. configurable [RSMM05]. configuration [CHE05b]. conflicting [HBK06]. consistency [JH03, Jun03]. consumer [SM03]. content [vdWBSV06]. context [Pot09]. Continuous [MPM09, Ben05, MR00]. control [Can04, Dal06d]. controlling [Ebe98]. controls [BDIM06]. Conversion [Cha96]. cooperation [ADG^{+}04]. cooperative [EG07]. Coordinating [RHV00]. coordination [ADG^{+}04, CN03, EC03, MLC^{+}08, ORM03]. core [RH01]. Correction [Ano05]. Cost [Mad08a, MC09]. COTS [MY06, MRE08, YB07]. COTS-based [YB07]. Creating [HSBR03, RSB05]. CRIM [BE96]. criteria [CWZ06]. Critical [NWZ06, BHM01, HN07]. critique [AG04]. culture [SS07]. current [KSK00]. customers [MAR08]. customizable [GKP09, ŞCVB05]. Cycle [CC96, Sin96].

DaimlerChrysler [SS01]. data

[MWE^{+}09, SCR05, APK06]. database [FD98]. databases [SBvV06]. Date [Cha96]. de-motivators [BHO07]. decentralized [TD08]. decision [AHI^{+}05, RCHSD00]. decisions [BMM07, YB07]. decomposition [BWRC05]. defect [WGJ^{+}08]. defects [GWY^{+}09]. defined [KK00, LGA07]. Defining [RBW07]. delivery [DC08]. demands [SM03]. Demotivators [NBK08]. Denmark [JC09]. departments [SGB06]. dependent [HBK06]. Deploying [KK00, AEH^{+}08]. derivation [RSMM05]. describe [XOZ^{+}07]. Description [EJ07]. Design [CC96, Dal07c, dJH96, BWRC05, DN07, FD98, GG08, GKP09, GGB03, GRSZ07, MV07, Nor09, PhD09b, Sch07, WMR04]. Design-based [CC96]. detection [Mad08b]. developer [BHJ06]. developers [DD03, Dal04a]. Developing [BJ07]. Development [AMK08, BM09, Ket07, Lar08, SFF^{+}06, TBD08, Ano09, ASAB02, APK06, dOBDVW06, Ben05, BG04, BKD08, Byg05, CLM^{+}08, CS00, CZS06, CHA06, Dal06d, Dam03, DSL08, DC08, EJ07, FG06, FGGM04, Ger03, GGB03, GWY^{+}09, HH00, Hen01, LK06, LMSW08, MBL07, Mad08a, MT97, MR00, MAR08, MD06, ÓOÅF09, ORM03, PL03, PAE03, RM00, SA07, oSWR07, Šmi06, SG06, SF08, Tho98, WMR04, WGHS01, WMM05, YB07, YP07, dAB07, CR09]. DEVS [CBK06]. different [Lin98, HBK06]. dimensional [BHO07]. direction [KSK00]. disciplines [Dal05a]. discrete [AEPR08, MR00, Pad02]. discrete-event [AEPR08]. dispersed [LMC03]. distance [WND08]. distributed [ADG^{+}04, Mad08a, PDL08, Zop09]. distribution [Šmi06, WGJ^{+}08]. do [Dal08]. Documentation [MRS06]. Does [WND08, Kau98]. domains [Pot09]. draw [SBvV06]. driven [dOBDVW06, LK06, NE05, SP08]. during [BG04]. Dynamic [BG04]. Dynamics [PAER07, AEP08, Byg05, DLH02, Häb04, HH00, MT00, RCHSD00].

e-commerce [DC08]. Early [Miy96, CM02]. Economics [DP08]. Edinburgh [PR05]. editor [RPW08, Dum98]. Editorial [Ano00, Ano01, GL08, Pau96, PS96a, PS96d, PS96b, PS96c, PS97a, PS97b, PS97c, PS97d, PS98a, PS98b, PS98c, PST00, PST01a, PST01b, PST01c, PST02a, PST02b, RS00]. Editorials [R06, AMK08, Ano98, BM06, BM07, BM08, BM09, Bos05a, CGA07, DR03a, DR03b, DR03c, Dal04a, Dal04b, DR04, DR05, Dal05a, Dal05b, Dal05c, Dal06a, Dal06c, Dal06b, Dal06d, Dal06e, Dal07a, Dal07b, Dal07c, Dal07d, Dal08, Dal09a, Dal09b, DR09, Dam03, DSL08, Din06, GRSZ07, KB03, MP06, PR05, PhD09a, PhD09b, RPW08, RSB05, Rou07, SR02, SFF^{+}06, WS04, Wer07, vOW05]. education [dOBDVW06]. educational [NvdH05]. effect [ASKM09]. Effective [DP96, MTK^{+}08]. effectiveness [Mad08b, YP07]. Effects [LMSW08]. efficiency [DLW06, YP07]. efficient [Tho98]. effort [Haa07, WGJ^{+}08]. electronic [MD06]. elicitation [CZS06, TM01, TT06]. embedded [KFS06, Pot07]. Empirical [ASKM09, BDW97, EG96, BKD08, CLM^{+}08, HRB02, Lin98, NWZ06, NBK08]. enacted [JPH07]. enactment [FGGM04]. encountered [TOK96]. end [DR09]. engineer [BHO07]. Engineering [R06, BE96, Lar08, Rom96, Wal03, BJ07, CZS06, KB03, LVL^{+}06, LM04, Lep06, MTK^{+}08, NKM08, NvdH05, Pot07, Pot09, RMCGSR07, RTM09, Rol09, SM03, vdWBSV06, EF97, JNH96]. enhancements [dOBDVW06]. enhancing [Dal06d]. Enterprise [Ket07]. enterprises [Kau98]. Entry [KM98]. Envelopment [APK06]. Environment [Ban96, ADG^{+}04, Dal06e, JSZ97]. Eoin [Ano09]. era [DR09]. eridging [MTK^{+}08]. Errata [Ano09]. ESI [Doi97, Sat97]. Establishing [JN96]. Establishment [SP08, Wie97]. estimates [MWE^{+}09]. Estimating [WGJ^{+}08]. estimation [ANB07]. Europe [BWHW06, Din06]. European [Mon96, And97, BD96, BM07, MŠWM06, Sat97, SMH96]. EuroSPI [BM06, Din06]. evaluate [oSWR07, WMR04]. Evaluating [JH03, Jun03, MTCC06, RFoSS08, GO05]. Evaluation [ANB07, CWZ06, EG96, BKD08, Lep06, MY06, RM00, Sai03, WMM05]. evaluations [CM02]. event [AEPR08]. Evolution [Rou03, BG04, CLRW00, OM09, RS02, Sal09, SCR05, SCFR06, WH02]. evolve [BG04]. Evolving [Gar97, NS07, KH08]. Examining [PhD09b]. excellence [DD03, Dal09a]. exist [Dal06a]. expanding [DR05]. Experience [DP96, JC09, Miy96, Ver97, ABS05, dOBDVW06, BM06, Col07, Hou06, LGM04, Moi98, SS01, Qin07]. experience-based [Moi98]. Experiences [AEH^{+}08, DDA05, Oja09, BM07, BM09, DKC00, Hen01, Hol97, IPB01, RGdASE08, SGB06]. experiment [ABEL97, BIM00, JSZ97, Kil97, Tho98]. Experimental [EF97]. Experimenting [Dal09b]. experiments [WMR04]. exploring [Ano09, ÓOÅF09]. Extending [Ket07]. external [BDIM06]. extreme [CM06].

facilitating [Hou06]. factor [MERO08]. factors [GdAVS08, MY06, NWZ06, SM98, Zop09]. Failures [DT02]. families [AHI^{+}05, Bos05b, KFS06, RSMM05]. fault [ASKM09, Mad08b]. Faults [DLW06]. Faults-slip-through [DLW06]. fear [CR08]. feature [CHE05a, CHE05b]. Features [CFQ08]. Fifth [Mon96]. financial [DHP^{+}07, ST07]. Findings [JHGEE01]. fine [ABEL97]. fine-grained [ABEL97]. first [MTCC06]. fit [RSB05]. Fitzgerald [Ano09]. fixing [WGJ^{+}08]. flexibility [GRSZ07, RBW07]. flexible [DN07, NE05, Sch07, SWG^{+}07]. foci [HBK06]. form [SWG^{+}07]. Formal [dJH96]. formalism [CBK06]. Formalizing [CHE05a]. formedness [TNT^{+}05]. forward [DR05]. foundation [EC03]. Four [FR05]. Framework [Miy96, TT06, dJH96, And04, Ano05, BDVD05, BJ07, DN07, FvGB03, GO05, Haa07, LY01, RRT04, SR07, BCQ02]. frameworks [Vak97]. France [MLN^{+}03]. Fraunhofer [EF97]. Free [SFF^{+}06, CHA06, CWZ06, SG06]. Free/ Open [SFF^{+}06, CWZ06, SG06]. functional [BWRC05]. Future [Dal07a, CG98, DR05, KSK00].

game [dOBDVW06, NvdH05]. gap [SM03]. gaps [KB03]. GENESIS [ADG^{+}04]. geographical [Šmi06]. geographically [LMC03]. Global [AMK08, BM09, Dam03, DSL08, PAE03, Šmi06, Ano09, EC03, Ger03, MLC^{+}08, Mad08a, MŠWM06, MTK^{+}08, ÓOÅF09, PL03, oSWR07, SB06, WH02, CR09, MŠ08]. globally [PDL08, Zop09]. GNOME [Ger03]. goals [HBK06]. going [Hum07]. good [SSV97]. grained [ABEL97]. granularity [Lin98]. grid [BH02]. group [BE96]. growing [Dam03]. growth [MŠWM06]. Guest [R06, AMK08, BM06, BM07, Bos05a, CGA07, Din06, GL08, GRSZ07, MP06, Pau96, PR05, RPW08, RSB05, Rou07, SFF^{+}06, Dum98]. guide [MD06]. Guided [FGGM04]. Guidelines [MTK^{+}08, BDR96, CJ08].

harmonized [MV07]. Harmonizing [RT07]. Hawaii [Pau96]. HCI [CN03]. headed [DSL08]. Helena [Ano09]. help [RBW07]. Heuristic [WSR05]. high [BHJ06]. Holmström [Ano09]. Holon [BCQ02]. Houses [SMH96]. Human [MERO08, KB03]. hungary [BIM00]. hybrid [CBK06, DI01, MBL07, WMR04].

IDEAL [CR04]. Identification [MR02]. identifying [Sof05]. II [DHP^{+}07]. IMMoS [PR02]. Impact [Mad08b, CR08, EC03, MTCC06, RFoSS08, Sof05, YP07]. impede [GdAVS08]. Implementation [CR09, AGV04, NWZ06, vdWBSV06]. implementations [RGdASE08]. Implementing [HRB02, Oja08, And04]. ImprovAbility [CJ08]. improve [Dal05c]. Improvement [Ano05, BCQ02, Dal07a, Dal07d, JN96, Li07, SMH96, Tom96, ABM07, AFT02, BH02, Ben05, BWHW06, BIM00, BM07, BM08, BDR96, CVO^{+}04, CFS01, DR03a, DR03b, DR03c, Dal04a, Dal04b, DR04, DR05, Dal05c, Dal09a, DKC00, Din06, FSC^{+}08, Gar97, GdAVS08, HRB02, Har05, HNOS09, Hou06, Hum07, HRW97, HW04, IPB01, JPH07, KM05, KM98, Kau98, KSK00, LT98, LK97, LGM04, LSK^{+}06, MPM09, MOC06, MERO08, Moi98, MR06, MRW09, MLN^{+}03, NWZ06, NBK08, PB98, RH01, RA05, Ram06, Rou07, RRT04, SA07, SGB06, SSV97, SP08, Sta00, SM98, TM97, Tho98, Vak97, Ano98]. improvements [SR07]. Improving [LR97, YB07, BDIM06, Haa07, PYM^{+}06, RTM09, WGHS01]. improvisation [SF08]. Inc. [Miy96]. incorporating [MY06]. Incremental [Dal04b, MBL07]. indicators [MV07]. industrial [BM06, BM07, Col07, MR06, NKM08, Oja08, PB98, RA05]. industries [MŠWM06]. industry [BSN07, CRa08, CLM^{+}08, DKC00, FFL^{+}07, ST07, Wie97]. informal [DD03]. informática [FSC^{+}08]. Information [CHA06, BKD08, Byg05, JS05, Sal09]. Informatique [ESM96]. infrastructure [JS05]. initial [WH02]. initiative [KM98]. initiatives [Moi98]. injection [WGJ^{+}08]. Innovation [JC09, BM07, BM09]. inquiry [HNOS09]. Inspection [TOK96, Har05, LMC03]. Institute [EF97, And97, BD96, JNH96, Rom96]. institutions [DHP^{+}07]. instructional [DDT01]. Intangible [HW04]. Integrated [Li07, MRH^{+}09, RTM09, RM00, Ano05, PR02, RRT04]. integrating [GGB03]. integration [GL08, Hou06]. Intensive [Mad08a, Ako09, GdAVS08]. Intention [NE05]. Intention-driven [NE05]. inter [PL03]. inter-organizational [PL03]. interaction [KB03]. interactive [BDVD05]. internal [JH03, Jun03]. International [Pau96, PR05, Tom96, WS04, MP06, Rou03, EG96, Sin96]. interoperability [GL08]. Interoperable [Lar08]. interorganizational [Sch07]. Introducing [And04, MT97]. Introduction [Ano98, DR03a, DR03b, DR03c, DR04, DR05, Dum98, RPW08, SSV97]. intrusive [TGC^{+}08]. invariants [RBW07]. investigating [ASKM09]. investigation [NBK08]. Isabel [Ano05]. ISBN [Ben05]. ISD [Lep06]. ISO [Sin96, Col07, Gar97, Jun03, Oul96, Rou03, RT07, SER^{+}97, Sta00, SMH96, Ben05]. ISO/IEC [Sin96, Col07, Gar97, Jun03, Rou03, RT07, SER^{+}97, Sta00]. Isolation [GG08]. Issue [Ano98, Pau96, WS04, Wer07, AMK08, BM06, DR03a, DR03b, DR04, KB03, MP06, PR05, SR02, vOW05]. issues [AGV04, BWHW06, LT98, PhD09a]. iterative [GWY^{+}09, SA07].

J [Ano09]. Japan [Tom96]. Jazz [SF08]. JIL [XOZ^{+}07].

key [Dal06b]. Knowledge

[BAM98, MR02, Sal09, XOZ^{+}07, FG06].

lack [MŠ08]. landscaping [Gru00]. large [And04, AEH^{+}08, BM08, GGdASA06, HH00, KL08, MPM09]. large-scale [And04, KL08]. Latvia [Šmi06]. lean [MAR08]. learn [Ost07]. learned [PAE03, SGB06]. Learning [DT02, BAM98, MOC06, ME07, MERO08, SF08]. learnt [BM08]. legacy [CM06, KH08, Sne98]. legal [BRMG09, RMCGSR07]. lessons [BM08, Dal05b, PAE03, SGB06, SBvV06]. Letter [Hum96]. Level [AGV04, CFQ08, Wal03, Dal06c, FR05, RTM09]. levels [BRMG09, Lin98]. Life [Sin96]. lifecycle [CM02, LMSW08]. Lightweight [MC09, CRa08, JSZ97]. Line [vOW05, CGC06, FR05, Pot09, RA05]. lines [AKM^{+}09, SB05]. Linking [DKC00]. Little [XOZ^{+}07]. Little-JIL [XOZ^{+}07]. local [MWE^{+}09]. long [CLRW00, SR07]. long-term [CLRW00, SR07]. looking [DR05]. Louis [MP06]. low [CJ08, WCM^{+}07]. low-maturity [CJ08]. Low-overhead [WCM^{+}07].

MA [vOW05]. maintaining [RSB05]. making [AHI^{+}05]. manage [RHV00]. Management [FG06, MRS06, dOBDVW06, BDVD05, BT98, BAM98, Bos05a, Dal06d, DHP^{+}07, GWY^{+}09, Haa07, Kil97, LGM04, NNR96, PhD09a, RMCGSR07, SB05, vdWBSV06, MR02, Wal03]. manager [BMM07]. Managing [Ako09, ADG^{+}04, Byg05, CM06, Dal08, DDT01, HRG00, Mid97, Moi98, Qin07, Sal09, ŞCVB05]. market [LK06]. market-driven [LK06]. marketplace [MŠWM06]. marriage [Sec09]. matching [TT06]. matrix [Ric01]. matter [WND08]. mature [NV98]. Maturing [SS01]. Maturity [KCF^{+}96, RM00, Vak97, BHJ06, BRMG09, CJ08, Gar97]. May [MP06, PR05, WS04]. measurable [LGA07]. Measurement [Dum98, ST07, APK06, BDR96, KFS06, PR02, Wie97]. measurement-based [BDR96]. measurements [CGA07]. measures [BMM07, CHA06, JH03, Jun03]. Measuring [Sne98, DLW06, Lin98]. mechanism [FFL^{+}07]. mechanisms [FR05]. medium [MT97, MD06, SR07, WMM05]. medium-sized [MD06, SR07, WMM05]. Mercedes [Ano05]. meta [FR05]. meta-level [FR05]. metamodel [DLH02]. metaphor [SF08]. Method [Lar08, Rol09, R06, AEPR08, AR07, CRa08, CZS06, IPB01, SBvV06, Sch07, TGC^{+}08, WCM^{+}07, vdWBSV06]. Methodological [AGV04]. methodology [AG04, BKD08, EJ07, PR02, Sal09, SF08]. methodology-tailoring [BKD08]. Methods [dJH96, DN07, GO05, LK06, Lep06, Rol09]. metrics [ASKM09, AR07, FD98, Hol97, RHV00]. Miguel [Ano05]. MinimalEDoc [Sal09]. Minimally [TGC^{+}08]. minimize [Sai03]. mismatch [MRE08]. Missouri [MP06]. Model [dOBDVW06, CC96, PAER07, RM00, AEP08, AR07, ASAB02, BKD08, Can04, CR04, DI01, GKP09, HSBR03, IPB01, MR00, NKM08, NNR96, Pad02, RMCGSR07, Ric01, Li07, Vak97]. Model-driven [dOBDVW06]. Modeling [KCF^{+}96, PR05, WS04, AFT02, BWRC05, Bro04, DLH02, FGGM04, JS05, MT00, Mad08a, MP06, NvdH05, NE05, OM09, RCHSD00, Sca00, SR02, TD08, Ver97]. modell [BWHW06]. Modelling [CLRW00, ABEL97, KH08, PR02, WGHS01]. models [And04, CHE05a, CHE05b, Gar97, Häb04, MR00, RHV00, RS02, Sof05, WSR05, WH02]. moderating [SG06]. modest [AFT02]. Monitoring [Lin98, AR07]. Montreal [ESM96]. motivation [BHJ06]. motivators [BHO07]. moving [Dal09a]. MPS.BR [MRW09]. Multi [LM04, BHO07, ORM03]. Multi-project [LM04]. multi-site [ORM03]. multilevel [CHE05b]. multiple [MŠ08]. multiple-case [MŠ08]. mutual [Byg05].

Nanda [Ben05]. navigate [ABM07]. NEC [KSK00]. need [Dal08]. Never [Dal05a]. News [And97, BD96, BAM98, BDW97, Doi97, ESM96, EF97, JNH96, Mon96, Rom96, Sat97, Ver97, Wie97]. next [Dal05c, Dal06c]. No [Ben05]. Non [CFQ08, Col07]. non-software [Col07]. Non-technical [CFQ08]. Norway [LK97]. novel [ASAB02]. nuclear [HN07].

Ó [Ano09]. object

[ASKM09, BDW97, MT97].

object-oriented [ASKM09, BDW97, MT97]. objectives [DC08]. observations [TGG08]. Oerlikon [LT98]. off [Kau98]. Olsson [Ano09]. one [Dal07c, Šmi06]. ongoing [Dam03]. OOSPICE [HSBR03]. open [ASAB02, CLM^{+}08, CHA06, Ger03, NT06, PYM^{+}06, SCR05, SCFR06]. open-source [PYM^{+}06]. Operational [DHP^{+}07, AEP08, AG04]. opportunities [Dam03, TBD08]. optimal [Pad06]. optimization [WSR05]. Optimized [MRE08]. organisation [Byg05]. organisational [Dal09a]. organisations [MOC06, SR07, TGM04, WMM05]. organization [MPM09]. Organizational [CS00, KL08, PL03, SM98, Ano98]. organizations [BM08, BAM98, CJ08, GdAVS08, Pot09, Qin07, YP07]. oriented [ASKM09, BDW97, LMSW08, LSK^{+}06, MT97, Sch07]. outsourcers [MLC^{+}08]. outsourcing [BDIM06, RCHSD00, SB06]. overhead [WCM^{+}07]. owners [TD08].

PA [CFQ08]. pages [Ben05]. pair [Mad08b, MTCC06]. Papers [KL08, Pau96, TBD08, BM06, Rou07]. paradigm [MT97]. paradigms [Gar97]. participation [dAB07]. path [CG98]. pattern [GKP09, Har05]. pattern-based [GKP09]. pay [Kau98]. PCRs [KM05]. PDTR [Gar97, SER^{+}97]. hardware [RM00]. IEC [Sin96, Col07, Gar97, Jun03, Rou03, RT07, SER^{+}97, Sta00]. Open [SFF^{+}06, CWZ06, SG06]. ProSim [Wer07]. people [ABS05, LT98, Vak97]. Performance [BMM07, PYM^{+}06, SP08, SG06]. Performing [AEP08]. personal [Hum07]. Perspective [Dal07d, TM01]. Perspectives [AMK08]. pervasive [Dal04b]. Phase [JN96, JHGEE01]. phased [SSV97]. Pioneering [BIM00]. Planning [JN96, AEPR08, AEP08, MR06, XOZ^{+}07]. Plans [PAER07]. plants [HN07]. policies [Pad02]. Portland [WS04]. postmortem [SBvV06]. postmortems [DDA05]. potential [BWHW06]. power [HN07]. Practical [BDR96, Sai03, AFT02, CR04, GGdASA06, Moi98, PB98, RGdASE08]. Practice [Dum98, R06, CN03, DR03c, DR05, Dal06a, KB03, MLC^{+}08, PAE03, RA05, Sat97, SSV97, vdWBSV06, Ano05]. practices [BSN07, PL03, ST07]. Practitioner [BH02]. prediction [Lin98]. preliminary [Kil97]. Preparing [TGC^{+}08]. Press [Ben05]. principles [BJ07, MAR08]. priorities [HBK06]. prioritising [Dal05c]. prioritization [LK06, YB07]. prioritizations [HBK06]. Prioritizing [KM05]. Probabilistic [DP08]. problem [Hou06, Šmi06]. problems [TOK96]. procedure [RM00]. Process [Ano98, Ano05, BCQ02, BM08, Cha96, Dal07a, Dal07d, DP96, DP08, DHP^{+}07, JS05, Li07, MR02, MRS06, Miy96, Mon96, PR05, SMH96, TOK96, Tom96, TD08, Vak97, WS04, XOZ^{+}07, AFT02, AR07, And04, ASAB02, AGJ^{+}00, ABEL97, AEH^{+}08, AKM^{+}09, AG04, BH02, Ban96, BWHW06, BIM00, BM07, BJ07, Bos05a, BDR96, Bro04, BRMG09, Can04, CVO^{+}04, CFS01, CLRW00, CGC06, CBK06, CS00, CZS06, DR03a, DR03b, DR03c, DR04, DR05, Dal05a, Dal05b, Dal05c, Dal06c, Dal09a, DR09, DLW06, DPV97, DLH02, DKC00, Din06, DI01, EJ07, FSC^{+}08, GKP09, GGB03, GO05, Gru00, HRB02, Har05, HH00, HNOS09, Hou06, Hum07, HRW97, HW04, JSZ97, Jun01, Jun03, KK00, KM05, KM98, Kau98, KB03, Kil97, KSK00, KH08, LT98, LK97, LGM04, LGA07]. process [LY01, Lin98, LSK^{+}06, MT00, MV07, MPM09, MR00, MTCC06, MOC06, MY06, MD06, Moi98, MR06, MRW09, MP06, NKM08, NvdH05, NWZ06, NBK08, OBM05, OM09, PhD09b, Pot07, Pot09, Qin07, RH01, RBW07, RMCGSR07, Ric01, Rou07, RRT04, SA07, SGB06, SSV97, Sca00, SR02, Sec09, SWG^{+}07, Sof05, Sta00, SM98, TM01, TM97, Tho98, TNT^{+}05, Ver97, WMR04, WSR05, Wal03, WMM05, WCM^{+}07, YB07, vWVS06, Dal06b]. process-centred [JSZ97]. Processes [BHM01, SFF^{+}06, ABS05, ADG^{+}04, BWRC05, BG04, Car07, Col07, CGA07, CG98, Dal08, Dal09b, DN07, EBS96, EG07, FGGM04, GdAVS08, GRSZ07, JSZ97, MBL07, Mad08a, NT06, Nor09, Ost07, PYM^{+}06, RSB05, RM00, Sch07, TGG08, WH02, WGHS01, R06, Sin96]. Product [Ket07, Kil97, Pot09, vOW05, AHI^{+}05, Bos05b, CGC06, FR05, LK06, RSMM05, RA05, SB05]. productivity [APK06, Dal06d]. products [CWZ06, Oja08, Oja09]. PROFES [AMK08]. professional [SS07]. program [LR97, MRW09]. programme [LGM04, SP08]. programming [CM06, Mad08b, MTCC06]. programs [IPB01, Wie97]. Project [BCQ02, Ket07, dOBDVW06, BMM07, CWZ06, DDA05, Ger03, LM04, NNR96, PDL08, Pad02, SBvV06, SG06, TT06, XOZ^{+}07]. projects [ABM07, Ako09, ASAB02, Byg05, Haa07, Häb04, KL08, LY01, ME07, NT06, PL03, Pad06, RHV00, Šmi06, Zop09]. promote [dAB07]. proneness [ASKM09]. properties [TNT^{+}05]. proposed [MŠWM06]. ProSim [MP06, PR05, WS04]. Prospective [EG96]. Providing [MAR08]. Pursuing [CFS01].

QFD [LSK^{+}06]. Qualification [HN07, FFL^{+}07]. Qualitative [RS02, SCR05]. Quality [Ben05, LK97, Tom96, AR07, BDW97, CM02, LR97, Mid97, MLN^{+}03, PYM^{+}06, SB06, SS07]. Quantitative [GWY^{+}09]. quest [Ebe98].

R&D [BE96]. Radiotherapy [BCQ02]. Ramos [Ano05]. rapid [Bro04]. Rating [Jun01, EBS96, JH03]. Rationale [OM09, BWRC05]. re [AEPR08, NKM08]. re-engineering [NKM08]. re-planning [AEPR08]. real [And04, Hol97, MC09]. real-world [And04]. really [FSC^{+}08]. Recherche [ESM96]. recommendations [Sai03]. redesign [Sca00]. Reduction [CC96]. reflecting [R06]. reflection [Qin07]. reflections [Dal05b]. regard [JH03]. Regular [KL08, TBD08]. regulatory [BDIM06]. relationship [FvGB03]. Release [MR06, PAER07, AEP08]. releases [AEPR08]. relevance [HBK06]. reliability [SER^{+}97]. Reorganizing [TOK96]. replanning [AEP08]. replicated [ASKM09]. Report [Ban96, DP96, Mon96, ABS05]. reported [RGdASE08]. reporting [Hou06]. reports [DDA05]. repositories [Hen01]. repository [Sat97]. requests [KM05]. Requirements [Nor09, SSV97, R06, Ako09, BDIM06, CS00, CZS06, HBK06, LVL^{+}06, LK06, Pot07, Pot09, RFoSS08, TT06, R06]. requires [JPH07]. Research [Dal06a, Miy96, SR07, Wal03, BWHW06, BDW97, MLN^{+}03, Rom96]. resolution [MRE08]. resolved [TOK96]. resources [MERO08]. Results [Bro04, AEH^{+}08, Kil97, SMH96]. retrospective [CG98]. reusability [Sne98]. reuse [GG08, HSBR03, NS07, PB98, Ram06]. Review [Ben05]. right [JPH07]. rigorous [AGJ^{+}00]. risk [AEP08, CVO^{+}04, DHP^{+}07, LGM04, TGC^{+}08, YB07]. risk-based [YB07]. risks [Sai03]. ROI [FSC^{+}08]. Role [Dal07d, FG06, SGB06, SG06, dAB07]. roles [BH02, BWRC05, JPH07]. room [Pen00]. Ruiz [Ano05].

safety [BHM01, HN07, MRH^{+}09]. safety-critical [HN07]. SAM [CFQ08]. scale [And04, HH00, JH03, KL08]. scaling [BHO07]. Scenario [AHI^{+}05]. Scenario-based [AHI^{+}05]. schedule [WGJ^{+}08]. scheduling [Pad02, Pad06]. Science [Pau96]. Scope [Sof05, DR05]. Scoping [AKM^{+}09]. Scotland [PR05]. scrum [PDL08, SF08]. SE [CN03]. Section [BAM98]. Sections [Jun03]. security [LVL^{+}06]. SEI [KCF^{+}96]. Selected [Rou07, BM06, Din06]. selecting [Dal05c]. selection [MRE08]. Self [SS96, DD03, ŞCVB05]. self-assessment [DD03]. Self-Assessments [SS96]. self-customizable [ŞCVB05]. sensitivity [WMR04]. Separation [CN03, EC03]. September [vOW05]. Sercomp [Sch07]. service [BDIM06, LMSW08, Sch07]. service-oriented [LMSW08]. services [NV98, Rol09]. Shifts [HBK06]. Short [Hum96, PS96b, PS96c]. simple [WH02]. simplification [DPV97]. Simulating [HH00, WH02]. Simulation [PR05, PAER07, WS04, AEPR08, AEP08, Ano05, ASAB02, CBK06, CS00, DI01, FGGM04, GKP09, Hou06, MTCC06, MP06, NvdH05, Pad02, PR02, RFoSS08, RS02, RRT04, Sca00, SR02, oSWR07, SCR05, SCFR06, Sta00, WMR04, WSR05]. simulation-based [Ano05, RRT04]. Simulations [CM02]. simulator [BDVD05, CGC06]. SISOS [MBL07]. site [ORM03]. Situational [R06, CZS06, LVL^{+}06, Lep06, vdWBSV06]. size [KFS06, MT97]. sized [MD06, SR07, WMM05]. skills [RTM09]. SLA [RMCGSR07]. slip [DLW06]. small [BD00, BM08, HRG00, Kau98, LY01, NNR96, Pen00, Ric01, SR07, TGC^{+}08, Tho98, TGM04, Vak97, WMM05, YP07, vWVS06]. small- [SR07, WMM05]. SMEs [Kil97]. social [CS00]. sociotechnical [Tax05]. SoftPM [XOZ^{+}07]. Software [AMK08, And97, Ano98, Ano05, BHJ06, BD96, Ban96, BCQ02, BWHW06, BM09, Bos05a, CR09, CC96, CG98, Dal05b, Dal06c, Dal06b, Dal07a, DR09, Din06, Dum98, EF97, Gru00, Hol97, Hum07, JNH96, Kau98, Ket07, KSK00, Li07, Mad08a, Miy96, MŠ08, Mon96, MLN^{+}03, NKM08, NvdH05, OBM05, PR05, PAER07, Ric01, Rom96, Rou07, SFF^{+}06, SB06, Sin96, Sta00, SS96, SMH96, Tom96, WS04, vOW05, AFT02, Ako09, AEPR08, AR07, And04, Ano09, ASAB02, AGJ^{+}00, AKM^{+}09, AG04, APK06, ADG^{+}04, BH02, BHO07, dOBDVW06, BHM01, Ben05, BG04, BDVD05, BJ07, BAM98, Bos05b, Bro04, CRa08, Can04, CFS01, CLRW00, CGC06, CLM^{+}08, CBK06, CS00, Col07, CZS06, CHA06, CWZ06, CGA07, DD03, DR03a, DR03b, DR03c, DR04, DR05, Dal05c, Dal06d]. software [Dal08, Dam03, DSL08, DLH02, DKC00, DDT01, DI01, DC08, EC03, FFL^{+}07, FG06, FGGM04, FSC^{+}08, FvGB03, FD98, FR05, GdAVS08, GKP09, Ger03, GGB03, Haa07, Häb04, HRB02, Har05, HH00, HNOS09, HRW97, HW04, JSZ97, KM05, KB03, KL08, KFS06, Kil97, LK97, LGM04, LM04, LGA07, LK06, LMSW08, LR97, LSK^{+}06, MLC^{+}08, MT00, MT97, MPM09, MR00, MŠWM06, MAR08, MTCC06, Mid97, MTK^{+}08, MD06, Moi98, MRW09, MP06, NS07, NWZ06, NBK08, ÓOÅF09, OM09, ORM03, PL03, Pad02, Pad06, PYM^{+}06, PAE03, Qin07, RSMM05, RHV00, RH01, RS02, RMCGSR07, RCHSD00, RRT04, RM00, Sai03, SA07, Sat97, ST07, Sca00, SR02, SB05, SM03, oSWR07, Šmi06, SCR05, Sne98, SM98, SG06, SF08, TGC^{+}08, TM97, TGG08, Ver97, WMR04, WSR05]. software [WH02, WMM05, WCM^{+}07, YP07, dAB07, vWVS06, Ano05, BE96, MR02, Miy96, RM00, Wal03]. Software-Intensive [Mad08a, Ako09]. software/hardware [RM00]. Some [TGG08, RH01]. Source [SFF^{+}06, ASAB02, CLM^{+}08, CHA06, CWZ06, Ger03, NT06, PYM^{+}06, SS07, SCR05, SCFR06, SG06]. SPACE [CVO^{+}04]. Spanish [GGdASA06]. Special [Ano98, BM06, KB03, MP06, Pau96, PR05, SR02, WS04, Wer07, vOW05, AMK08]. specialization [CHE05a, CHE05b]. Specifying [dJH96]. SPI [ABM07, BHO07, BD00, BM09, DC08, GGdASA06, HRG00, MC09, Pot07, Ric01]. SPICE [EG96, CVO^{+}04, EBS96, FFL^{+}07, JHGEE01, Jun01, JH03, MRH^{+}09, Miy96, Pot07, Rou07, SR07, TGM04]. SPICE-based [FFL^{+}07, Jun01]. SPIdex [FSC^{+}08]. SPIP [Wer07]. spiral [AR07, NNR96]. SPLC [vOW05]. SPW [Wer07]. SPW/ProSim [Wer07]. St [MP06]. Stability [PAER07]. stage [SG06]. Staged [Bos05b, CHE05b]. Standard [Sin96, vWVS06, AEH^{+}08, Rou03, EG96]. standardization [Mid97]. Standards [Jun03, DR03a, Hen01, LVL^{+}06, SR07]. started [Hum07]. state [MLC^{+}08]. state-of-the-practice [MLC^{+}08]. static [LR97]. status [KSK00]. step [Dal07c]. still [WND08]. stochastic [Can04]. stories [DDA05]. strategies [DKC00, KM98, MERO08, NS07, oSWR07]. strategy [HH00, PB98]. Streamline [TBD08]. Strengthening [BRMG09]. structured [CR09]. structures [Häb04]. studie [RH01]. Studies [DPV97, AR07, MT00]. Study [BCQ02, GGdASA06, ANB07, ASKM09, BHJ06, CLRW00, CLM^{+}08, Ger03, HRB02, JC09, Lin98, MLC^{+}08, MŠ08, MR06, NKM08, NNR96, NWZ06, Oja08, Ost07, PDL08, Pad06, RGdASE08, SCR05, Vak97, WMR04, Wal03, WH02]. Success [JPH07, SM98, Zop09, BDIM06, CHA06, Dal06b, Dal07b, NWZ06]. Successful [JN96, MRW09, Wie97]. Suitability [LK06]. suites [Mad08b]. supplier [FFL^{+}07]. suppliers [MLC^{+}08]. Support [SS96, FGGM04, GRSZ07, HRW97, LMC03, RSB05, RCHSD00, Sch07, SWG^{+}07, Sta00, XOZ^{+}07, dAB07]. Supporters [FG06]. Supporting [CFQ08, Dal06d, BMM07, Dal04a, Dal05a, JSZ97]. survey [BSN07]. Surveys [SMH96]. sustaining [Dal09a]. SW [KK00]. symbiosis [JPH07]. SynQuest [SS96]. System [PAER07, RCHSD00, AEP08, BKD08, DLH02, Häb04, KFS06, MT00, RTM09, RM00, Sal09, vdWBSV06]. system-applications [vdWBSV06]. Systematic [SB05]. Systems [GL08, Lar08, Mad08a, Pau96, And04, BHM01, BG04, Byg05, CM06, CHA06, FR05, GGB03, GRSZ07, HN07, HH00, KH08, LM04, ME07, Pot07, RSB05, Sne98, ŞCVB05].

tailored [TGC^{+}08]. tailoring [AEH^{+}08, BKD08]. target [Dal09a]. task [oSWR07]. taxonomies [MY06]. teaching [CN03]. Team [MRS06, BD00, JSZ97, Mad08a, Pen00, Tho98]. Team-based [MRS06]. Teams [Dal07a, CR08, EC03, LMC03, YP07, dAB07, MŠ08]. technical [CFQ08, Ebe98]. technique [AFT02, BH02]. Techniques [PYM^{+}06, TT06]. technologies [BT98]. technology [Byg05, Mon96]. telecom [MLN^{+}03]. Ten [GdAVS08]. term [CLRW00, SR07]. test [Can04, DLW06, Mad08b, MTCC06]. test-first [MTCC06]. their [BG04, CHE05a, MLC^{+}08, RSMM05]. Theory [Dum98, CN03, CHA06]. thinking [Dal09b, LM04]. thoroughness [Mad08b]. three [Kil97]. Time [CC96, Dal07c, EC03]. Tool [HRW97, LMC03, SS96, ANB07, ORM03, WSR05]. tools [RFoSS08]. Topological [TNT^{+}05]. Toro [Ano05]. TR [Jun03]. Traditional [TBD08]. trials [CVO^{+}04, JHGEE01]. TRISO [Li07]. TRISO-Model [Li07]. trust [MŠ08]. TSP [RGdASE08]. turbulence [YP07]. Turning [Hen01]. TVO [HN07]. Two [SMH96, Sec09].

ubiquitous [Ost07]. UK [Tom96]. unadjusted [KFS06]. Understanding [MŠ08, Sca00, SFF^{+}06, CR08]. unexplored [Ano09, ÓOÅF09]. unit [Mad08b]. units [LGA07]. universal [Ako09]. USA [MP06]. Usability [NT06, FvGB03, GGB03, SM03]. use [MD06]. user [GGB03, TM01]. user-centered [GGB03]. Using [APK06, BHO07, DC08, PDL08, SM03, oSWR07, WMR04, AEPR08, AEP08, BH02, BJ07, CBK06, CM06, Col07, HH00, LGM04, LM04, LSK^{+}06, MR00, Miy96, Pot07, RFoSS08, SR07, Sca00, SCR05]. utilizing [Jun01].

V [BWHW06]. V&V [WSR05]. V-modell [BWHW06]. validation [AG04, GdAVS08, KFS06, vdWBSV06]. value [BJ07, MAR08, Oja08, Oja09]. value-based [BJ07]. variability [AHI^{+}05, Bos05a, RA05, ŞCVB05]. various [Pot09]. vendor [MY06]. Verification [AG04, GdAVS08]. versus [DDA05]. very [BD00, CM02, Kau98]. via [SMH96, YB07]. View [TM01, Hum07]. View-based [TM01]. viewpoint [HBK06]. viewpoint-dependent [HBK06]. Virtual [CR08, Pen00, SB06]. Vivek [Ben05]. volume [DR03c, DR05].

way [MŠWM06]. ways [Dal09b]. weapon [Dal04b]. Web [Car07, JS05, Pen00, vdWBSV06]. Web-based [Pen00, vdWBSV06]. well [DR03b, TNT^{+}05]. well-formedness [TNT^{+}05]. where [DSL08, Hum07]. within [MŠWM06]. without [MWE^{+}09]. work [Dal05a]. workflow [BT98, NE05, WGHS01]. Workflows [EG07]. Working [Dal06e]. Workshop [Ban96, Mon96, PR05, WS04, MP06]. workshops [CZS06]. world [And04, Hol97]. worlds [Sec09]. worldwide [BM08]. worth [FSC^{+}08].

XP [BSN07]. XT [BWHW06]. xv [Ben05].

Year [Cha96]. yields [HBK06].

References

Arlow:1997:FGP

[ABEL97] Jim Arlow, Sergio Bandinelli, Wolfgang Emmerich, and Luigi Lavazza. A fine-grained process modelling experiment at British Airways. Software Process: Improvement and Practice, 3(2):105–131, June 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Aaen:2007:SAH

[ABM07] Ivan Aaen, Anna Börjesson, and Lars Mathiassen. SPI agility: How to navigate improvement projects. Software Process: Improvement and Practice, 12(3):267–281, May 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Andersson:2005:APB

[ABS05] Tomas Andersson, Ilia Bider, and Rogier Svensson. Aligning people to business processes experience report. Software Process: Improvement and Practice, 10(4):403–413, October 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Aversano:2004:MCC

[ADG^{+}04] Lerina Aversano, Andrea De Lucia, Matteo Gaeta, Pierluigi Ritrovato, Silvio Stefanucci, and Maria Luisa Villani. Managing coordination and cooperation in distributed software processes: the GENESIS environment. Software Process: Improvement and Practice, 9(4):239–263, October 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Armbrust:2008:ERT

[AEH^{+}08] Ove Armbrust, Jan Ebell, Ulrike Hammerschall, Jürgen Münch, and Daniela Thoma. Experiences and results from tailoring and deploying a large process standard in a company. Software Process: Improvement and Practice, 13 (4):301–309, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Al-Emran:2008:POR

[AEP08] Ahmed Al-Emran and Dietmar Pfahl. Performing operational release planning, replanning and risk analysis using a system dynamics simulation model. Software Process: Improvement and Practice, 13 (3):265–279, May 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Al-Emran:2008:MRP

[AEPR08] Ahmed Al-Emran, Dietmar Pfahl, and Günther Ruhe. A method for re-planning of software releases using discreteevent simulation. Software Process: Improvement and Practice, 13(1):19–33, January 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ahonen:2002:MPS

[AFT02] Jarmo J. Ahonen, Marko Forsell, and Sanna-Kaisa Taskinen. A modest but practical software process modeling technique for software process improvement. Software Process: Improvement and Practice, 7(1):33–44, March 2002.

CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Arthur:2004:VVO

[AG04] James D. Arthur and Markus K. Gröner. Verification and validation of operational software: a process and methodology critique. Software Process: Improvement and Practice, 9 (3):157–171, July 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ares:2000:MRC

[AGJ^{+}00] Juan Ares, Rafael García, Natalia Juristo, Marta López, and Ana M. Moreno. A more rigorous and comprehensive approach to software process assessment. Software Process: Improvement and Practice, 5 (1):3–30, March 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Antoniol:2004:MIC

[AGV04] Giuliano Antoniol, Sara Gradara, and Gabriele Venturi. Methodological issues in a CMM Level 4 implementation. Software Process: Improvement and Practice, 9(1):33–50, January 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

America:2005:SBD

[AHI^{+}05] Pierre America, Dieter Hammer, Mugurel T. Ionita, Henk Obbink, and Eelco Rommes.

Scenario-based decision making for architectural variability in product families. Software Process: Improvement and Practice, 10(2):171–187, April 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Armbrust:2009:SSP

[AKM^{+}09] Ove Armbrust, Masafumi Katahira, Yuko Miyamoto, Jürgen Münch, Haruka Nakao, and Alexis Ocampo. Scoping software process lines. Software Process: Improvement and Practice, 14(3):181–197, May 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Akoumianakis:2009:MUA

[Ako09] Demosthenes Akoumianakis.

Managing universal accessibility requirements in softwareintensive projects. Software Process: Improvement and Practice, 14(1):3–29, January 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Abrahamsson:2008:GEP

[AMK08] Pekka Abrahamsson, Jürgen Münch, and Pasi Kuvaja. Guest editorials: Perspectives on global software development: special issue on PRO-FES 2007. Software Process: Improvement and Practice, 13 (3):213–215, May 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Abran:2007:EBB

[ANB07] Alain Abran, Iphigénie Ndiaye, and Pierre Bourque. Evaluation of a black-box estimation tool: a case study. Software Process: Improvement and Practice, 12(2):199–218, March 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Andres:1997:NES

[And97] Ana Andrés. News: European Software Institute. Software Process: Improvement and Practice, 3(3):189–190, September 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Andreou:2004:IPF

[And04] Andreas S. Andreou. Introducing a process framework for implementing models of large-scale real-world systems in software. Software Process: Improvement and Practice, 9 (3):133–155, July 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Anonymous:1998:EIS

[Ano98] Anonymous. Editorials: Introduction to special issue on ‘Organizational Change in Software Process Improvement’. Software Process: Improvement and Practice, 4(4):185, December 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Anonymous:2000:E

[Ano00] Anonymous. Editorial. Software Process: Improvement and Practice, 5(4):211, December 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Anonymous:2001:E

[Ano01] Anonymous. Editorial. Software Process: Improvement and Practice, 6(1):1, March 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Anonymous:2005:CIF

[Ano05]  Anonymous.  Correction:

“An integrated framework for simulation-based Software Process Improvement”, by Mercedes Ruiz, Isabel Ramos, Miguel Toro, Software Process Improvement and Practice 2004; 9:81–93. Software Process: Improvement and Practice, 10(3):355, July 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic). See [RRT04].

Anonymous:2009:EBG

[Ano09] Anonymous. Errata: “Benefits of global software development: exploring the unexplored”, Eoin Ó Conchúir, Helena Holmström Olsson, Pär J. Ågerfalk, Brian Fitzgerald. Software Process: Improvement and Practice, 14(5): 301, September 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic). See [ÓOÅF09].

Asmild:2006:UDE

[APK06] Mette Asmild, Joseph C. Paradi, and Atin Kulkarni. Using Data Envelopment Analysis in software development productivity measurement. Software Process: Improvement and Practice, 11(6): 561–572, November 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Andersson:2007:SPM

[AR07] Carina Andersson and Per Runeson. A spiral process model for case studies on software quality monitoring — method and metrics. Software Process: Improvement and Practice, 12(2):125–140, March 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Antoniades:2002:NSM

[ASAB02] I. P. Antoniades, I. Stamelos, L. Angelis, and G. L. Bleris. A novel simulation model for the development process of open source software projects. Software Process: Improvement and Practice, 7(3–4):173–188, September/December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Aggarwal:2009:EAI

[ASKM09] K. K. Aggarwal, Yogesh Singh, Arvinder Kaur, and Ruchika Malhotra. Empirical analysis for investigating the effect of object-oriented metrics on fault proneness: a replicated case study. Software Process: Improvement and Practice, 14 (1):39–62, January 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bomarius:1998:NSK

[BAM98] Frank Bomarius, Klaus-Dieter Althoff, and Wolfgang Müller. News section: Knowledge management for learning software organizations. Software Process: Improvement and Practice, 4(2):89–93, June 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bandinelli:1996:RWS

[Ban96] Sergio Bandinelli. Report on the workshop of software process environment architecture. Software Process: Improvement and Practice, 2(1):54–72, March 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bell:2002:HFS

[BCQ02] G. A. Bell, M. A. Cooper, and S. Qureshi. The Holon Framework and software process improvement: A radiotherapy project case study. Software Process: Improvement and Practice, 7(2):57–70, June 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bandinelli:1996:NES

[BD96] Sergio Bandinelli and Amor Domínguez. News: European Software Institute. Software Process: Improvement and Practice, 2(4):311–312, December 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Batista:2000:SVS

[BD00] J. Batista and A. Dias De Figueiredo. SPI in a very small team: a case with CMM. Software Process: Improvement and Practice, 5(4):243– 250, December 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Biro:2006:CBS

[BDIM06] Miklós Biró, Csilla Deák, János Ivanyos, and Richard Messnarz. From compliance to business success: improving outsourcing service controls by adopting external regulatory requirements. Software Process: Improvement and Practice, 11(3):239–249, May 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Briand:1996:PGM

[BDR96] Lionel C. Briand, Christiane M. Differding, and H. Dieter Rombach. Practical guidelines for measurement-based process improvement. Software Process: Improvement and Practice, 2(4):253–280, December 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Birkholzer:2005:ISM

[BDVD05] Thomas Birkhölzer, Christoph Dickmann, Jürgen Vaupel, and Laura Dantas. An interactive software management simulator based on the CMMI framework. Software Process: Improvement and Practice, 10 (3):327–340, July 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Briand:1997:NER

[BDW97] Lionel Briand, John Daly, and Juergen Wuest. News: Empirical research in object-oriented quality. Software Process: Improvement and Practice, 3(4): 247–251, December 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Briand:1996:SER

[BE96] Lionel Briand and Khaled El Emam. Software Engineering R&D group at CRIM. Software Process: Improvement and Practice, 2(1):73–74, March 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Benediktsson:2005:BRI

[Ben05] Professor Oddur Benediktsson.

Book review: ISO 9001:2000: Achieving compliance and continuous improvement in software development companies. Vivek Nanda, ASQ Quality Press, 2003. No. of pages: xv + 271. ISBN 0-87389-594-0. Software Process: Improvement and Practice, 10(1):97–98, January 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Beydeda:2004:DES

[BG04] Sami Beydeda and Volker Gruhn. Dynamic evolution of software processes to evolve software systems during their development. Software Process: Improvement and Practice, 9(4):229–238, October 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Baddoo:2002:PRS

[BH02] Nathan Baddoo and Tracy Hall. Practitioner roles in software process improvement: an analysis using grid technique. Software Process: Improvement and Practice, 7(1): 17–31, March 2002. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Baddoo:2006:SDM

[BHJ06] Nathan Baddoo, Tracy Hall, and Dorota Jagielska. Software developer motivation in a high maturity company: a case study. Software Process: Improvement and Practice, 11 (3):219–228, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Benediktsson:2001:PSS

[BHM01] O. Benediktsson, R. B. Hunter, and A. D. McGettrick. Processes for software in safety critical systems. Software Process: Improvement and Practice, 6(1):47–62, March 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Baddoo:2007:UMD

[BHO07] Nathan Baddoo, Tracy Hall, and Ciaran O’Keeffe. Using multi dimensional scaling to analyse software engineers’ de-motivators for SPI. Software Process: Improvement and Practice, 12(6):511– 522, November 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Biro:2000:PPI

[BIM00] Miklós Biró, János Ivanyos, and Richard Messnarz. Pioneering process improvement experiment in hungary. Software Process: Improvement and Practice, 5(4):213–229, December 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Boehm:2007:DPF

[BJ07] Barry Boehm and Apurva Jain. Developing a process framework using principles of value-based software engineering. Software Process: Improvement and Practice, 12(5): 377–385, September 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Burns:2008:EEM

[BKD08] Timothy Burns, Robb Klashner, and Fadi Deek. An empirical evaluation of a methodology-tailoring information system development model. Software Process: Improvement and Practice, 13(5): 387–395, September 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Biro:2006:GES

[BM06] Miklós Biró and Richard Messnarz. Guest editorials: Special issue with selected industrial experience papers of EuroSPI^{2} 2005. Software Process: Improvement and Practice, 11(3): 215–218, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Biro:2007:GEE

[BM07] Miklós Biró and Richard Messnarz. Guest editorials: European industrial experiences in process improvement and innovation. Software Process: Improvement and Practice, 12(6): 507–509, November 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Biro:2008:EPI

[BM08] Miklós Biró and Richard Messnarz. Editorials: Process improvement lessons learnt in small and large organizations worldwide. Software Process:

Improvement and Practice, 13 (4):297–299, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Biro:2009:ESE

[BM09] Miklós Biró and Richard Messnarz. Editorials: SPI experiences and innovation for global software development. Software Process: Improvement and Practice, 14(5):243– 245, September 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bertolino:2007:PMS

[BMM07] Antonia Bertolino, Eda Marchetti, and Raffaela Mirandola. Performance measures for supporting project manager decisions. Software Process: Improvement and Practice, 12(2): 141–164, March 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bosch:2005:GES

[Bos05a] Jan Bosch. Guest editorials: Software variability: process and management. Software Process: Improvement and Practice, 10(1):3–5, January 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bosch:2005:SAS

[Bos05b] Jan Bosch. Staged adoption of software product families. Software Process: Improvement and Practice, 10(2):

125–142, April 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Buglione:2009:SML

[BRMG09] Luigi Buglione, Ricardo J. Rejas-Muslera, and Juan José Cuadrado Gallego. Strengthening maturity levels by a legal assurance process. Software Process: Improvement and Practice, 14(6):305–314, November 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Brooks:2004:RRB

[Bro04] Andrew Brooks. Results of rapid bottom-up software process modeling. Software Process: Improvement and Practice, 9(4):265–278, October 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bowers:2007:AXP

[BSN07] Alexia N. Bowers, Raghvinder S. Sangwan, and Colin J. Neill. Adoption of XP practices in the industry — a survey. Software Process: Improvement and Practice, 12(3): 283–294, May 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bolcer:1998:AWM

[BT98] Gregory Alan Bolcer and Richard N. Taylor. Advanced workflow management technologies. Software Process: Improvement and Practice, 4 (3):125–171, September 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Biffl:2006:SPI

[BWHW06] Stefan Biffl, Dietmar Winkler, Reinhard Höhn, and Herbert Wetzel. Software process improvement in Europe: potential of the new V-modell XT and research issues. Software Process: Improvement and Practice, 11(3):229–238, May 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Balabko:2005:CDR

[BWRC05] Pavel Balabko, Alain Wegmann, Alain Ruppen, and Nicolas Clément. Capturing design rationale with functional decomposition of roles in business processes modeling. Software Process: Improvement and Practice, 10(4): 379–392, October 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Bygstad:2005:MDM

[Byg05] Bendik Bygstad. Managing the dynamics of mutual adaptation of technology and organisation in information systems development projects. Software Process: Improvement and Practice, 10(3):341–353, July 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Cangussu:2004:STP

[Can04] João W. Cangussu. A software test process stochastic control model based on CMM characterization. Software Process: Improvement and Practice, 9 (2):55–66, April 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Cardoso:2007:CAB

[Car07] Jorge Cardoso. Complexity analysis of BPEL Web processes. Software Process: Improvement and Practice, 12(1): 35–49, January 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Choi:2006:AHS

[CBK06] KeungSik Choi, Doo-Hwan Bae, and TagGon Kim. An approach to a hybrid software process simulation using the DEVS formalism. Software Process: Improvement and Practice, 11(4):373–383, July 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Collier:1996:DBM

[CC96] Ken W. Collier and James S. Collofello. A design-based model for the reduction of software cycle time. Software Process: Improvement and Practice, 2(3):167–179, September 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Carvallo:2008:SCL

[CFQ08] Juan Pablo Carvallo, Xavier Franch, and Carme Quer. Supporting CMMI Level 2 SAM PA with non-technical features catalogues. Software Process: Improvement and Practice, 13 (2):171–182, March 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Cattaneo:2001:PCS

[CFS01] F. Cattaneo, A. Fuggetta, and D. Sciuto. Pursuing coherence in software process assessment and improvement. Software Process: Improvement and Practice, 6(1):3–22, March 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Cugola:1998:SPR

[CG98] Gianpaolo Cugola and Carlo Ghezzi. Software processes: a retrospective and a path to the future. Software Process: Improvement and Practice, 4(3): 101–123, September 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Cuadrado-Gallego:2007:GEA

[CGA07] Juan J. Cuadrado-Gallego and Alain Abran. Guest editorials: Advances in measurements for software processes assessment. Software Process: Improvement and Practice, 12(2): 121–123, March 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Chen:2006:SPL

[CGC06] Yu Chen, Gerald C. Gannod, and James S. Collofello. A software product line process simulator. Software Process: Improvement and Practice, 11 (4):385–409, July 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Chavan:1996:YDC

[Cha96] Subhash Chavan. The year 2000 date conversion process. Software Process: Improvement and Practice, 2(2):111– 122, June 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Crowston:2006:ISS

[CHA06] Kevin Crowston, James Howison, and Hala Annabi. Information systems success in free and open source software development: theory and measures. Software Process: Improvement and Practice, 11(2): 123–148, March 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Czarnecki:2005:FCB

[CHE05a] Krzysztof Czarnecki, Simon Helsen, and Ulrich Eisenecker. Formalizing cardinality-based feature models and their specialization. Software Process: Improvement and Practice, 10 (1):7–29, January 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Czarnecki:2005:SCT

[CHE05b] Krzysztof Czarnecki, Simon Helsen, and Ulrich Eisenecker. Staged configuration through specialization and multilevel configuration of feature models. Software Process: Improvement and Practice, 10(2): 143–169, April 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Christiansen:2008:IGL

[CJ08] Mads Christiansen and Jørn Johansen. ImprovAbility guidelines for low-maturity organizations. Software Process: Improvement and Practice, 13 (4):319–325, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Chen:2008:ESS

[CLM^{+}08] Weibing Chen, Jingyue Li, Jianqiang Ma, Reidar Conradi, Junzhong Ji, and Chunnian Liu. An empirical study on software development with open source components in the Chinese software industry. Software Process: Improvement and Practice, 13(1): 89–100, January 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Chatters:2000:MSE

[CLRW00] B. W. Chatters, M. M.

Lehman, J. F. Ramil, and P. Wernick. Modelling a software evolution process: a long-term case study. Software Process: Improvement and Practice, 5(2–3):91–102, June/September 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Chiang:2002:SVE

[CM02] Eliza Chiang and Tim Menzies. Simulations for very early lifecycle quality evaluations. Software Process: Improvement and Practice, 7(3–4):141–159, September/December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Coleman:2006:MCL

[CM06] Gerry Coleman and Martin McAnallen. Managing the challenges of legacy systems using extreme programming. Software Process: Improvement and Practice, 11(3):269– 275, May 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Clemmensen:2003:STC

[CN03] Torkil Clemmensen and Jacob Nørbjerg. Separation in theory, coordination in practice — teaching HCI and SE. Software Process: Improvement and Practice, 8(2):99– 110, April 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Coletta:2007:IEA

[Col07]  Antonio Coletta.  An in-

dustrial experience in assessing the capability of nonsoftware processes using ISO/ IEC 15504. Software Process: Improvement and Practice, 12 (4):315–319, July 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Casey:2004:PAI

[CR04] Valentine Casey and Ita Richardson. A practical application of the IDEAL model. Software Process: Improvement and Practice, 9(3):123– 132, July 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Casey:2008:VTU

[CR08] Valentine Casey and Ita Richardson. Virtual teams: understanding the impact of fear. Software Process: Improvement and Practice, 13(6): 511–526, November 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Casey:2009:IGS

[CR09] Valentine Casey and Ita Richardson. Implementation of Global Software Development: a structured approach. Software Process: Improvement and Practice, 14(5):247– 262, September 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Caffery:2008:AAL

[CRa08] Fergal Mc Caffery, Ita Richardson, and Peter a Moller. Automotive–adept: a lightweight assessment method for the automotive software industry. Software Process: Improvement and Practice, 13 (4):345–353, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Christie:2000:OSS

[CS00] Alan M. Christie and Mary Jo Staley. Organizational and social simulation of a software requirements development process. Software Process: Improvement and Practice, 5(2– 3):103–110, June/September 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Cass:2004:SST

[CVO^{+}04] Ann Cass, Christian Völcker, Rafik Ouared, Alec Dorling, Lothar Winzer, and Juan María Carranza. SPICE for SPACE trials, risk analysis, and process improvement. Software Process: Improvement and Practice, 9(1): 13–21, January 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Cruz:2006:ECF

[CWZ06] David Cruz, Thomas Wieland, and Alexander Ziegler. Evaluation criteria for free/open source software products based on project analysis. Software Process: Improvement and Practice, 11(2):107–122, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Coulin:2006:SME

[CZS06] Chad Coulin, Didar Zowghi, and Abd-El-Kader Sahraoui. A situational method engineering approach to requirements elicitation workshops in the software development process. Software Process: Improvement and Practice, 11(5): 451–464, September 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

deAraujo:2007:RCS

[dAB07] Renata Mendes de Araujo and Marcos R. S. Borges. The role of collaborative support to promote participation and commitment in software development teams. Software Process: Improvement and Practice, 12 (3):229–246, May 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2004:ECI

[Dal04a] Darren Dalcher. Editorials:

Capability improvement: supporting developers. Software Process: Improvement and Practice, 9(4):215–216, October 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2004:EIC

[Dal04b] Darren Dalcher. Editorials: Incremental capability improvement: the new weapon in addressing the all-pervasive complexity. Software Process: Improvement and Practice, 9(3):

121–122, July 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2005:ENA

[Dal05a] Darren Dalcher. Editorials: Never alone: process work and supporting disciplines. Software Process: Improvement and Practice, 10(3):249, July 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2005:ESP

[Dal05b] Darren Dalcher. Editorials: Software process: lessons and reflections. Software Process: Improvement and Practice, 10 (2):99–100, April 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2005:EWI

[Dal05c] Darren Dalcher. Editorials: What to improve next? selecting and prioritising in software process improvement. Software Process: Improvement and Practice, 10(4):357–358, October 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2006:ERP

[Dal06a] Darren Dalcher. Editorials: Research and practice can coexist. Software Process: Improvement and Practice, 11 (3):213, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2006:ESPb

[Dal06b] Darren Dalcher. Editorials:

Software Process: the key to success. Software Process: Improvement and Practice, 11 (2):93, March 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2006:ESPa

[Dal06c] Darren Dalcher. Editorials:

Software process: the next level. Software Process: Improvement and Practice, 11 (1):1–2, January 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2006:ESS

[Dal06d] Darren Dalcher. Editorials:

Supporting software development: enhancing productivity, management and control. Software Process: Improvement and Practice, 11(6):557– 559, November 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2006:EWC

[Dal06e] Darren Dalcher. Editorials:

Working in a complex environment. Software Process: Improvement and Practice, 11 (4):337, July 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2007:EAS

[Dal07a] Darren Dalcher. Editorials:

Agility and software teams: The future of software process improvement. Software Process: Improvement and Practice, 12(3):219–221, May 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2007:ECS

[Dal07b] Darren Dalcher. Editorials: Celebrating success. Software Process: Improvement and Practice, 12(6):505–506, November 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2007:EDC

[Dal07c] Darren Dalcher. Editorials: Design for change: one step at a time. Software Process: Improvement and Practice, 12 (1):1–2, January 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2007:ERP

[Dal07d] Professor Darren Dalcher. Editorials: The role of perspective in process improvement. Software Process: Improvement and Practice, 12(2):119–120, March 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2008:EMS

[Dal08] Darren Dalcher. Editorials: Managing software processes: do we need new approaches? Software Process: Improvement and Practice, 13(5):383– 385, September 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2009:EAS

[Dal09a] Darren Dalcher. Editorials:

Attaining and sustaining organisational excellence: process improvement as a moving target. Software Process: Improvement and Practice, 14 (2):63–64, March 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2009:EEN

[Dal09b] Darren Dalcher. Editorials:

Experimenting with new ways of thinking about processes. Software Process: Improvement and Practice, 14(1):1– 2, January 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Damian:2003:EGS

[Dam03] Daniela Damian. Editorials:

Global software development: growing opportunities, ongoing challenges. Software Process: Improvement and Practice, 8(4):179–182, October 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Downey:2008:USA

[DC08] Fergal Downey and Gerry Coleman. Using SPI to achieve delivery objectives in e-commerce software development. Software Process: Improvement and Practice, 13 (4):327–333, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Daily:2003:TSE

[DD03] K. Daily and D. Dresner. Towards software excellence — informal self-assessment for software developers. Software Process: Improvement and Practice, 8(3):157–168, July 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Desouza:2005:ECP

[DDA05] Kevin C. Desouza, Torgeir Dingsøyr, and Yukika Awazu. Experiences with conducting project postmortems: reports versus stories. Software Process: Improvement and Practice, 10(2):203–215, April 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Demirors:2001:MIS

[DDT01] Onur Demirors, Elif Demirors, and Ayca Tarhan. Managing instructional software acquisition. Software Process: Improvement and Practice, 6(4): 189–203, December 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

DiRenzo:2007:ORM

[DHP^{+}07] B. Di Renzo, M. Hillairet, M. Picard, A. Rifaut, C. Bernard, D. Hagen, P. Maar, and D. Reinard. Operational risk management in financial institutions: Process assessment in concordance with Basel II. Software Process: Improvement and Practice, 12 (4):321–330, July 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Donzelli:2001:HSP

[DI01] Paolo Donzelli and Giuseppe Iazeolla. A hybrid software process simulation model. Software Process: Improvement and Practice, 6(2):97– 109, June 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dingsoyr:2006:GES

[Din06] Torgeir Dingsøyr. Guest editorials: Software process improvement in Europe: selected articles from EuroSPI 2004. Software Process: Improvement and Practice, 11 (1):3–5, January 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

dInverno:1996:FFS

[dJH96] Mark d’Inverno, G. R. Ribeiro Justo, and Paul Howells. A formal framework for specifying design methods. Software Process: Improvement and Practice, 2(3):181–196, September 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Debou:2000:LSP

[DKC00] Christophe Debou and Annie Kuntzmann-Combelles. Linking software process improvement to business strategies: experiences from industry. Software Process: Improvement and Practice, 5(1):55–64, March 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

DeOliveiraBarros:2002:SDM

[DLH02] Márcio De Oliveira Barros, Cláudia Maria Lima Werner, and Guilherme Horta Travassos. A system dynamics metamodel for software process modeling. Software Process: Improvement and Practice, 7(3–4):161–172, September/December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Damm:2006:FST

[DLW06] Lars-Ola Damm, Lars Lundberg, and Claes Wohlin. Faults-slip-through — a concept for measuring the efficiency of the test process. Software Process: Improvement and Practice, 11(1):47–59, January 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Daoudi:2007:BFM

[DN07] Feriel Daoudi and Selmin Nurcan. A benchmarking framework for methods to design flexible business processes.

Software Process: Improvement and Practice, 12(1):51– 63, January 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Barros:2006:MDG

[dOBDVW06] Márcio de O. Barros, Alexandre R. Dantas, Gustavo O. Veronese, and Cláudia M. L. Werner. Model-driven game development: experience and model enhancements in software project management education. Software Process: Improvement and Practice, 11 (4):411–421, July 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Doiz:1997:NEN

[Doi97] Izaskun Doiz. News: ESI news. Software Process: Improvement and Practice, 3(1): 62–63, March 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dandekar:1996:BEP

[DP96] Ashok Dandekar and Dewayne E. Perry. Barriers to effective process architecture — an experience report. Software Process: Improvement and Practice, 2(1):13–19, March 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Deissenboeck:2008:PAP

[DP08] Florian Deissenboeck and Markus Pizka. Probabilistic analysis of process economics. Software Process: Improvement and Practice, 13 (1):5–17, January 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dandekar:1997:SPS

[DPV97] Ashok Dandekar, Dewayne E. Perry, and Lawrence G. Votta. Studies in process simplification. Software Process: Improvement and Practice, 3(2): 87–104, June 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2003:EIIa

[DR03a] Darren Dalcher and David Raffo. Editorials: Introduction to issue 8:1: software process improvement standards. Software Process: Improvement and Practice, 8(1):3–4, January 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2003:EIIb

[DR03b] Darren Dalcher and David Raffo. Editorials: Introduction to issue 8:3: software process improvement is alive and well. Software Process: Improvement and Practice, 8 (3):133–134, July 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2003:EIS

[DR03c] Darren Dalcher and David Raffo. Editorials: Introduction to software process: improvement and practice, volume 8. Software Process: Improvement and Practice, 8 (1):1–2, January 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2004:EII

[DR04] Darren Dalcher and David Raffo. Editorials: Introduction to issue 9 : 1: software process improvement for all. Software Process: Improvement and Practice, 9(1):1–2, January 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2005:EIS

[DR05] Darren Dalcher and David Raffo. Editorials: Introduction to software process: improvement and practice, volume 10 looking forward to the future: expanding the scope of the process area. Software Process: Improvement and Practice, 10 (1):1–2, January 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2009:ESP

[DR09] Darren Dalcher and David Raffo. Editorials: Software process: the end of an era. Software Process: Improvement and Practice, 14(6):303– 304, November 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Damian:2008:EGS

[DSL08] Daniela Damian, Bikram Sengupta, and Filippo Lanubile.

Editorials: Global software development: where are we headed? Software Process: Improvement and Practice, 13 (6):473–475, November 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dalcher:2002:LF

[DT02] Darren Dalcher and Colin Tully. Learning from failures. Software Process: Improvement and Practice, 7(2):71–89, June 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Dumke:1998:GEI

[Dum98] Reiner Dumke. Guest Editor introduction: Software measurement in theory and practice. Software Process: Improvement and Practice, 4(1): 19, March 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ebert:1998:QTC

[Ebe98] Christof Ebert. The quest for technical controlling. Software Process: Improvement and Practice, 4(1):21–31, March 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

ElEmam:1996:AAR

[EBS96] Khaled El Emam, Lionel Briand, and Robert Smith. Assessor agreement in rating SPICE processes. Software Process: Improvement and Practice, 2(4):291–306, December 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Espinosa:2003:ITS

[EC03] J. Alberto Espinosa and Erran Carmel. The impact of time separation on coordination in global software teams: a conceptual foundation. Software Process: Improvement and Practice, 8(4): 249–266, October 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

ElEmam:1997:NFI

[EF97] Khaled El Emam and Pierfrancesco Fusaro. News: Fraunhofer Institute for Experimental Software Engineering. Software Process: Improvement and Practice, 3(1):59–61, March 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

ElEmam:1996:EEP

[EG96] Khaled El Emam and Dennis R. Goldenson. An empirical evaluation of the prospective International SPICE Standard. Software Process: Improvement and Practice, 2(2): 123–148, June 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Estublier:2007:WCP

[EG07] Jacky Estublier and Sergio Garcia. Workflows and cooperative processes. Software Process: Improvement and Practice, 12(5):415–427, September 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Eberlein:2007:DPD

[EJ07] Armin Eberlein and Li Jiang. Description of a process development methodology. Software Process: Improvement and Practice, 12(1):101–118, January 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

ElEmam:1996:NCR

[ESM96] Khaled El Emam, Barry Shostak, and Nazim H. Madhavji. News: Centre de recherche informatique de montreal. Software Process: Improvement and Practice, 2 (2):155–158, June 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Foltin:1998:ASM

[FD98] Erik Foltin and Reiner R. Dumke. Aspects of software metrics database design. Software Process: Improvement and Practice, 4(1):33–42, March 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Fabbrini:2007:SBS

[FFL^{+}07] Fabrizio Fabbrini, Mario Fusani, Giuseppe Lami, Edoardo Sivera, and Edoardo Sivera. A SPICE-based software supplier qualification mechanism in automotive industry. Software Process: Improvement and Practice, 12(6):523–528, November 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Feher:2006:RKM

[FG06] Péter Fehér and András Gábor. The role of Knowledge Management Supporters in software development companies. Software Process: Improvement and Practice, 11 (3):251–260, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Fernandez:2004:GSC

[FGGM04] Alejandro Fernández, Badie Garzaldeen, Ines Grützner, and Jürgen Münch. Guided support for collaborative modeling, enactment and simulation of software development processes. Software Process: Improvement and Practice, 9 (2):95–106, April 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Fritsch:2005:FMA

[FR05] Claudia Fritsch and Burkhardt Renz. Four mechanisms for adaptable systems: a metalevel approach to building a software product line. Software Process: Improvement and Practice, 10(2):103–124, April 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ferreira:2008:RSP

[FSC^{+}08] Analia Irigoyen Ferreiro Ferreira, Gleison Santos, Roberta Cerqueira, Mariano Montoni, Ahilton Barreto, Ana Regina Rocha, Andrea O. S. Barreto, and Reinaldo C. Silva Filho. ROI of software process improvement at BL informática: SPIdex is really worth it. Software Process: Improvement and Practice, 13 (4):311–318, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Folmer:2003:FCR

[FvGB03] Eelke Folmer, Jilles van Gurp, and Jan Bosch. A framework for capturing the relationship between usability and software architecture. Software Process: Improvement and Practice, 8 (2):67–87, April 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Garcia:1997:EIP

[Gar97] Suzanne M. Garcia. Evolving improvement paradigms: capability maturity models and ISO/IEC 15504 (PDTR). Software Process: Improvement and Practice, 3(1):47–58, March 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Garcia:2008:TFI

[GdAVS08] Javier García, Antonio de Amescua, Manuel Velasco, and Ana Sanz. Ten factors that impede improvement of verification and validation processes in software intensive organizations. Software Process: Improvement and Practice, 13 (4):335–343, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

German:2003:GPC

[Ger03]  Daniel M. German.  The

GNOME project: a case study of open source, global software development. Software Process: Improvement and Practice, 8(4):201–215, October 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Gamble:2008:IDR

[GG08] M. T. Gamble and R. F.

Gamble. Isolation in design reuse. Software Process: Improvement and Practice, 13(2): 145–156, March 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Goransson:2003:UDP

[GGB03] Bengt Göransson, Jan Gulliksen, and Inger Boivie. The usability design process — integrating user-centered systems design in the software development process. Software Process: Improvement and Practice, 8(2):111–131, April 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Guzman:2006:CSP

[GGdASA06] Javier García Guzmán, Román López-Cortijo García, Antonio de Amescua Seco, and Gonzalo Cuevas Agustín. Case study: a practical approach for SPI in large Spanish companies. Software Process: Improvement and Practice, 11(3): 261–268, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Garousi:2009:CPB

[GKP09] Vahid Garousi, Keyvan Khosrovian, and Dietmar Pfahl. A customizable pattern-based software process simulation model: design, calibration and application. Software Process: Improvement and Practice, 14 (3):165–180, May 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Gamble:2008:GES

[GL08] Rose Gamble and Grace A. Lewis. Guest editorial: Systems interoperability, integration, and composition. Software Process: Improvement and Practice, 13(2):111–112, March 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Green:2005:FCE

[GO05] Stewart Green and Martyn Ould. A framework for classifying and evaluating process architecture methods. Software Process: Improvement and Practice, 10(4):415–425, October 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Green:2007:GEB

[GRSZ07] Stewart Green, Gil Regev, Pnina Soffer, and Jelena Zdravkovic. Guest editorials: Business processes and support systems: design for flexibility. Software Process: Improvement and Practice, 12 (1):3–5, January 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Gruhn:2000:SPL

[Gru00]  Volker Gruhn.  Software

process landscaping. Software Process: Improvement and Practice, 5(2–3):111–120, June/September 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Gou:2009:QDM

[GWY^{+}09] Lang Gou, Qing Wang, Jun Yuan, Ye Yang, Mingshu Li, and Nan Jiang. Quantitative defects management in iterative development with BiDefect. Software Process: Improvement and Practice, 14 (4):227–241, July 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Haapio:2007:FIE

[Haa07] Topi Haapio. A framework for improving effort management in software projects. Software Process: Improvement and Practice, 12(6):549–558, November 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Haberlein:2004:CSS

[Häb04] Tobias Häberlein. Common structures in system dynamics models of software acquisition projects. Software Process: Improvement and Practice, 9(2):67–80, April 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Harjumaa:2005:PAS

[Har05] Lasse Harjumaa. A pattern approach to software inspection process improvement. Software Process: Improvement and Practice, 10(4):455–465, October 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Hoorn:2006:SFP

[HBK06] Johan F. Hoorn, Mark E. Breuker, and Evelien Kok. Shifts in foci and priorities. Different relevance of requirements to changing goals yields conflicting prioritizations and is viewpoint-dependent. Software Process: Improvement and Practice, 11(5):465–485, September 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Henninger:2001:TDS

[Hen01]  Scott Henninger.  Turning

development standards into repositories of experiences. Software Process: Improvement and Practice, 6(3):141– 155, September 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Henderson:2000:SPS

[HH00] Peter Henderson and Yvonne Howard. Simulating a process strategy for large scale software development using systems dynamics. Software Process: Improvement and Practice, 5(2–3):121–131, June/ September 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Halminen:2007:QSC

[HN07] Juha Halminen and Risto Nevalainen. Qualification of safety-critical systems in TVO nuclear power plants. Software Process: Improvement and Practice, 12(6):559–567, November 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Holmberg:2009:AIS

[HNOS09] Lena Holmberg, Agneta Nilsson, Helena Holmström Olsson, and Anna Börjesson Sandberg. Appreciative inquiry in software process improvement. Software Process: Improvement and Practice, 14(2): 107–125, March 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Holt:1997:SMR

[Hol97] John Holt. Software metrics — real world experiences. Software Process: Improvement and Practice, 3(3):155– 163, September 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Houston:2006:EFP

[Hou06] Dan Houston. An experience in facilitating process improvement with an integration problem reporting process simulation. Software Process: Improvement and Practice, 11 (4):361–371, July 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Hall:2002:ISP

[HRB02] Tracy Hall, Austen Rainer, and Nathan Baddoo. Implementing software process improvement: an empirical study. Software Process: Improvement and Practice, 7(1):3–15, March 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Horvat:2000:MCS

[HRG00] Romana Vajde Horvat, Ivan Rozman, and József Györkös. Managing the complexity of SPI in small companies. Software Process: Improvement and Practice, 5(1):45–54, March 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Hunter:1997:TSS

[HRW97] Robin Hunter, Gordon Robinson, and Ian Woodman. Tool support for software process assessment and improvement. Software Process: Improvement and Practice, 3(4):213– 223, December 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Henderson-Sellers:2003:COM

[HSBR03] B. Henderson-Sellers, J. Bohling, and T. Rout. Creating the OOSPICE model architecture — a case of reuse. Software Process: Improvement and Practice, 8(1):41–49, January 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Humphrey:1996:SCL

[Hum96] Watts S. Humphrey. Short communications: Letter. Software Process: Improvement and Practice, 2(1):3–4, March 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Humphrey:2007:SPI

[Hum07] Watts S. Humphrey. Software process improvement — a personal view: How it started and where it is going. Software Process: Improvement and Practice, 12(3):223–227, May 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Hyde:2004:IBC

[HW04] Kevin Hyde and David Wilson. Intangible benefits of CMM-based software process improvement. Software Process: Improvement and Practice, 9(4):217–228, October 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Isacsson:2001:ACB

[IPB01] Peter Isacsson, Gunnar Pedersen, and Stig Bang. Accelerating CMM-based improvement programs: the accelerator model and method with experiences. Software Process: Improvement and Practice, 6 (1):23–34, March 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Johansen:2009:EIC

[JC09] Jørn Johansen and Mads Christiansen. Experience with innovation checks: a case study with 46 companies in denmark. Software Process: Improvement and Practice, 14(5): 263–270, September 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jung:2003:ESR

[JH03] Ho-Won Jung and Robin Hunter. Evaluating the SPICE rating scale with regard to the internal consistency of capability measures. Software Process: Improvement and Practice, 8 (3):169–178, July 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jung:2001:FPS

[JHGEE01] Ho-Won Jung, Robin Hunter, Dennis R. Goldenson, and Khaled El-Emam. Findings from Phase 2 of the SPICE trials. Software Process: Improvement and Practice, 6(4): 205–242, December 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jones:1996:EPP

[JN96] Lawrence G. Jones and Linda M.

Northrop. The establishing phase: Planning for successful improvement. Software Process: Improvement and Practice, 2(1):51–53, March 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jones:1996:NSE

[JNH96] Lawrence G. Jones, Linda M.

Northrop, and James D. Hart. News: Software Engineering Institute. Software Process: Improvement and Practice, 2 (2):149–152, June 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Johansen:2007:SIR

[JPH07] Jørn Johansen and Jan Pries-Heje. Success with improvement — requires the right roles to be enacted — in symbiosis. Software Process: Improvement and Practice, 12(6):

529–539, November 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jensen:2005:PMA

[JS05] Chris Jensen and Walt Scacchi. Process modeling across the Web information infrastructure. Software Process: Improvement and Practice, 10 (3):255–272, July 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jahnke:1997:EBL

[JSZ97] Jens H. Jahnke, Wilhelm Schäfer, and Albert Zündorf. An experiment in building a lightweight process-centred environment supporting team software processes. Software Process: Improvement and Practice, 3(3):141–153, September 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jung:2001:RPA

[Jun01] Ho-Won Jung. Rating the process attribute utilizing AHP in SPICE-based process assessments. Software Process: Improvement and Practice, 6(2): 111–122, June 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Jung:2003:SSE

[Jun03] Ho-Won Jung. Standards sections: Evaluating the internal consistency of ISO/IEC TR 15504 process capability measures. Software Process: Improvement and Practice, 8(1): 5–26, January 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kautz:1998:SPI

[Kau98] Karlheinz Kautz. Software process improvement in very small enterprises: does it pay off? Software Process: Improvement and Practice, 4(4): 209–226, December 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kazman:2003:ESI

[KB03] Rick Kazman and Len Bass.

Editorials: Special issue on bridging the process and practice gaps between software engineering and human– computer interaction. Software Process: Improvement and Practice, 8(2):63–65, April 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Konrad:1996:CMM

[KCF^{+}96] Mike Konrad, Mary Beth Chrissis, Jack Ferguson, Suzanne Garcia, Bill Hefley, Dave Kitson, and Mark Paulk. Capability maturity modeling^{SM} at the SEI. Software Process: Improvement and Practice, 2 (1):21–34, March 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kettunen:2007:ESP

[Ket07] Petri Kettunen. Extending software project agility with new product development enterprise agility. Software Process: Improvement and Practice, 12(6):541–548, November 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kiebusch:2006:USM

[KFS06] Sebastian Kiebusch, Bogdan Franczyk, and Andreas Speck. An unadjusted size measurement of embedded software system families and its validation. Software Process: Improvement and Practice, 11 (4):435–446, July 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kotonya:2008:CBP

[KH08] Gerald Kotonya and John Hutchinson. A componentbased process for modelling and evolving legacy systems. Software Process: Improvement and Practice, 13(2):113– 125, March 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kilpi:1997:PMC

[Kil97] Tapani Kilpi. Product management challenge to software change process: preliminary results from three SMEs experiment. Software Process: Improvement and Practice, 3(3): 165–175, September 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kaltio:2000:DDS

[KK00] Timo Kaltio and Atte Kinnula. Deploying the defined SW process. Software Process: Improvement and Practice, 5 (1):65–83, March 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kettunen:2008:RPC

[KL08] Petri Kettunen and Maarit Laanti. Regular papers: Combining agile software projects and large-scale organizational agility. Software Process: Improvement and Practice, 13(2): 183–193, March 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kasse:1998:ESP

[KM98] Tim Kasse and Patricia A.

Mcquaid. Entry strategies into the process improvement initiative. Software Process: Improvement and Practice, 4 (2):73–88, June 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Kanungo:2005:PPC

[KM05] Shivraj Kanungo and Ipininder S. Monga. Prioritizing process change requests (PCRs) in software process improvement. Software Process: Improvement and Practice, 10(4): 441–453, October 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Komiyama:2000:SPA

[KSK00] Toshihiro Komiyama, Toshihiko Sunazuka, and Shinji Koyama. Software process assessment and improvement in NEC — current status and future direction. Software Process: Improvement and Practice, 5(1):31–43, March 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Larrucea:2008:MEA

[Lar08]  Xabier Larrucea.  Method

engineering approach for interoperable systems development. Software Process: Improvement and Practice, 13(2): 127–133, March 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Leppanen:2006:CEM

[Lep06] Mauri Leppänen. Conceptual evaluation of methods for engineering situational ISD methods. Software Process: Improvement and Practice, 11(5): 539–555, September 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Lee:2007:CPA

[LGA07] Seok-Won Lee, Robin A. Gandhi, and Gail-Joon Ahn. Certification process artifacts defined as measurable units for software assurance. Software Process: Improvement and Practice, 12(2):165–189, March 2007. CODEN SPIPFL.

ISSN 1077-4866 (print), 1099-1670 (electronic).

Lassudrie:2004:EUR

[LGM04] Claire Lassudrie and Gina Gullà-Menez. An experience in using risk management in a software process improvement programme. Software Process: Improvement and Practice, 9 (1):3–12, January 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Li:2007:TMN

[Li07] Mingshu Li. TRISO-Model:

A new approach to integrated software process assessment and improvement. Software Process: Improvement and Practice, 12(5):387– 398, September 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Lindvall:1998:MMC

[Lin98] Mikael Lindvall. Monitoring and measuring the changeprediction process at different granularity levels: an empirical study. Software Process: Improvement and Practice, 4 (1):3–10, March 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Larsen:1997:QAS

[LK97] Even by Larsen and Karlheinz Kautz. Quality assurance and software process improvement in Norway. Software Process: Improvement and Practice, 3 (2):71–86, June 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Lehtola:2006:SRP

[LK06] Laura Lehtola and Marjo Kauppinen. Suitability of requirements prioritization methods for market-driven software product development. Software Process: Improvement and Practice, 11(1):7– 19, January 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Lee:2004:MPS

[LM04] Bengee Lee and James Miller. Multi-project software engineering analysis using systems thinking. Software Process: Improvement and Practice, 9 (3):173–214, July 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Lanubile:2003:TSG

[LMC03] Filippo Lanubile, Teresa Mallardo, and Fabio Calefato. Tool support for geographically dispersed inspection teams. Software Process: Improvement and Practice, 8(4):217– 231, October 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Lewis:2008:ESO

[LMSW08] Grace A. Lewis, Edwin Morris, Soumya Simanta, and Lutz Wrage. Effects of serviceoriented architecture on software development lifecycle activities. Software Process: Improvement and Practice, 13(2): 135–144, March 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Lichter:1997:ISQ

[LR97] Horst Lichter and Gerhard Riedinger. Improving software quality by static program analysis. Software Process: Improvement and Practice, 3(4): 235–241, December 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Liu:2006:BOS

[LSK^{+}06] Xiaoqing (Frank) Liu, Yan Sun, Gautam Kane, Yuji Kyoya, and Kunio Noguchi. Business-oriented software process improvement based on CMM using QFD. Software Process: Improvement and Practice, 11(6):573–589, November 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Laporte:1998:API

[LT98] Claude Y. Laporte and Sylvie Trudel. Addressing the people issues of process improvement activities at Oerlikon Aerospace. Software Process: Improvement and Practice, 4 (4):187–198, December 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Laleau:2006:ASR

[LVL^{+}06] Régine Laleau, Sylvie Vignes, Yves Ledru, Michel Lemoine, Didier Bert, Véronique Donzeau-Gouge, Catherine Dubois, and Fabien Peureux. Adopting a situational requirements engineering approach for the analysis of civil aviation security standards. Software Process: Improvement and Practice, 11 (5):487–503, September 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Leung:2001:PFS

[LY01] Hareton K. N. Leung and Terence C. F. Yuen. A process framework for small projects. Software Process: Improvement and Practice, 6(2):67–83, June 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Madachy:2008:CMD

[Mad08a] Raymond J. Madachy. Cost modeling of distributed team processes for global development and software-intensive systems of systems. Software Process: Improvement and Practice, 13(1):51–61, January 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Madeyski:2008:IPP

[Mad08b] Lech Madeyski. Impact of pair programming on thoroughness and fault detection effectiveness of unit test suites. Software Process: Improvement and Practice, 13(3):281–295, May 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Mehta:2008:PVC

[MAR08] Merwan Mehta, David Anderson, and David Raffo. Providing value to customers in software development through lean principles. Software Process: Improvement and Practice, 13(1):101–109, January 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Madachy:2007:AHI

[MBL07] Raymond Madachy, Barry Boehm, and Jo Ann Lane. Assessing hybrid incremental processes for SISOS development. Software Process: Improvement and Practice, 12(5): 461–473, September 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

McCaffery:2009:LSA

[MC09] Fergal McCaffery and Gerry Coleman. Lightweight SPI assessments: what is the real cost? Software Process: Improvement and Practice, 14(5): 271–278, September 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Moe:2006:UEP

[MD06] Nils Brede Moe and Tore Dybå. The use of an electronic process guide in a mediumsized software development company. Software Process: Improvement and Practice, 11 (1):21–34, January 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Messnarz:2007:ABL

[ME07] Richard Messnarz and Damjan Ekert. Assessment-based learning systems — learning from best projects. Software Process: Improvement and Practice, 12(6):569–577, November 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Messnarz:2008:HRB

[MERO08] Richard Messnarz, Damjan Ekert, Michael Reiner, and Gearoid O’Suilleabhain. Human resources based improvement strategies — the learning factor. Software Process: Improvement and Practice, 13 (4):355–362, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Middleton:1997:MSQ

[Mid97] Peter Middleton. Managing software quality by standardization. Software Process: Improvement and Practice, 3(4): 201–212, December 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Miyoshi:1996:EES

[Miy96] Takeshige Miyoshi. Early experience with software process assessment using SPICE framework at Software Research Associates, Inc. Software Process: Improvement and Practice, 2 (3):211–235, September 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ma:2008:SPS

[MLC^{+}08] Jianqiang Ma, Jingyue Li, Weibing Chen, Reidar Conradi, Junzhong Ji, and Chunnian Liu. A state-of-the-practice study on communication and coordination between Chinese software suppliers and their global outsourcers. Software Process: Improvement and Practice, 13 (3):233–247, May 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Moreau:2003:SQI

[MLN^{+}03] Bernard Moreau, Claire Lassudrie, Bertrand Nicolas, Odile l’Homme, Cécile d’Anterroches, and Gilbert Le Gall. Software quality improvement in France telecom research center. Software Process: Improvement and Practice, 8(3): 135–144, July 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Messnarz:2006:PIL

[MOC06] Richard Messnarz, Gearoid O’Suilleabhain, and Ray Coughlan. From process improvement to learning organisations. Software Process: Improvement and Practice, 11 (3):287–294, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Moitra:1998:MCS

[Moi98] Deependra Moitra. Managing change for software process improvement initiatives: a practical experience-based approach. Software Process: Improvement and Practice, 4(4): 199–207, December 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Montangero:1996:NRF

[Mon96] Carlo Montangero. News: Report on the Fifth European Workshop on Software Process Technology. Software Process: Improvement and Practice, 2 (4):307–311, December 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Munch:2006:GES

[MP06] Jürgen Münch and Dietmar Pfahl. Guest editorials: Special issue on ProSim 2005, The 6th international workshop on software process simulation and modeling, St. Louis, Missouri, USA May 2005. Software Process: Improvement and Practice, 11(4):339–343, July 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Malheiros:2009:CPI

[MPM09] Viviane Malheiros, Fábio Rilston Paim, and Manoel Mendonça. Continuous process improvement at a large software organization. Software Process: Improvement and Practice, 14(2):65–83, March 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Martin:2000:MSD

[MR00] Robert H. Martin and David Raffo. A model of the software development process using both continuous and discrete models. Software Process: Improvement and Practice, 5(2–3):147–157, June/ September 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Meehan:2002:ISP

[MR02] Bridget Meehan and Ita Richardson. Identification of Software Process Knowledge Management. Software Process: Improvement and Practice, 7(2):47–55, June 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Momoh:2006:RPP

[MR06] Joseph Momoh and Günther Ruhe. Release planning process improvement — an industrial case study. Software Process: Improvement and Practice, 11(3):295–307, May 2006.

CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Mohamed:2008:OMR

[MRE08] Abdallah Mohamed, Guenther Ruhe, and Armin Eberlein. Optimized mismatch resolution for COTS selection. Software Process: Improvement and Practice, 13(2):157–169, March 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Messnarz:2009:IAS

[MRH^{+}09] Richard Messnarz, Hans-Leo Ross, Stephan Habel, Frank König, Abdelhadi Koundoussi, Jürgen Unterrreitmayer, and Damjan Ekert. Integrated automotive SPICE and safety assessments. Software Process: Improvement and Practice, 14 (5):279–288, September 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Meisinger:2006:TBP

[MRS06] Michael Meisinger, Andreas Rausch, and Marc Sihling. 4everedit — team-based process documentation management. Software Process: Improvement and Practice, 11(6): 627–642, November 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Montoni:2009:MBS

[MRW09] Mariano Angel Montoni, Ana Regina Rocha, and Kival Chaves Weber. MPS.BR: a successful program for software process improvement in Brazil. Software Process: Improvement and Practice, 14(5):289– 300, September 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Moe:2008:ULT

[MŠ08] Nils Brede Moe and Darja Šmite. Understanding a lack of trust in Global Software Teams: a multiple-case study. Software Process: Improvement and Practice, 13(3):217– 231, May 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

McCaffery:2006:PWE

[MŠWM06] Fergal McCaffery, Darja Šmite, F. G. Wilkie, and D. McFall. A proposed way for European software industries to achieve growth within the global marketplace. Software Process: Improvement and Practice, 11 (3):277–285, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Makris:1997:IOO

[MT97] V. A. Makris and K. C.

Thrampoulidis. Introducing the object-oriented paradigm in a medium size software development company. Software Process: Improvement and Practice, 3(1):35–46, March 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Madachy:2000:CSS

[MT00] Ray Madachy and Denton Tarbet. Case studies in software process modeling with system dynamics. Software Process: Improvement and Practice, 5(2–3):133–146, June/ September 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Melis:2006:EIT

[MTCC06] Marco Melis, Ivana Turnu, Alessandra Cau, and Giulio Concas. Evaluating the impact of test-first programming and pair programming through software process simulation. Software Process: Improvement and Practice, 11(4):345– 360, July 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Milewski:2008:GEE

[MTK^{+}08] Allen E. Milewski, Marilyn Tremaine, Felix Köbler, Richard Egan, Suling Zhang, and Patrick O’Sullivan. Guidelines for effective eridging in global software engineering. Software Process: Improvement and Practice, 13(6):477– 492, November 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Makinen:2007:HDP

[MV07] Timo Mäkinen and Timo Varkoi. A harmonized design for process assessment indicators. Software Process: Improvement and Practice, 12 (4):331–338, July 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Menzies:2009:AEL

[MWE^{+}09] Tim Menzies, Steve Williams, Oussama Elrawas, Daniel Baker, Barry Boehm, Jairus Hihn, Karen Lum, and Ray Madachy. Accurate estimates without local data? Software Process: Improvement and Practice, 14(4):213–225, July 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Miller:2006:CAP

[MY06] James Miller and H. C. Yeoh.

COTS acquisition process: incorporating business factors into COTS vendor evaluation taxonomies. Software Process: Improvement and Practice, 11 (6):601–626, November 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Niazi:2008:DSP

[NBK08] Mahmood Niazi, Muhammad Ali Babar, and Nolin Mark Katugampola. Demotivators of software process improvement: an empirical investigation. Software Process: Improvement and Practice, 13 (3):249–264, May 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Nurcan:2005:IDM

[NE05] Selmin Nurcan and Marc-Henri Edme. Intention-driven modeling for flexible workflow applications. Software Process: Improvement and Practice, 10(4):363–377, October 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Napier:2008:SPR

[NKM08] Nannette P. Napier, Jonathan Kim, and Lars Mathiassen. Software process re-engineering: a model and its application to an industrial case study. Software Process: Improvement and Practice, 13(5):451– 471, September 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Nelson:1996:ASM

[NNR96] Howard C. Nelson, Tom Nute, and David J. Rodjak. Applying the spiral model: a case study in small project management. Software Process: Improvement and Practice, 2(4): 239–251, December 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Nori:2009:RDP

[Nor09] Kesav V. Nori. Requirements for design of processes. Software Process: Improvement and Practice, 14(3):129–142, May 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Nedstam:2007:ESS

[NS07] Josef Nedstam and Mark Staples. Evolving strategies for software architecture and reuse. Software Process: Improvement and Practice, 12 (3):295–309, May 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Nichols:2006:UPO

[NT06] David M. Nichols and Michael B.

Twidale. Usability processes in open source projects. Software Process: Improvement and Practice, 11(2):149–162, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Niessink:1998:TMI

[NV98] Frank Niessink and Hans Van Vliet. Towards mature IT services. Software Process: Improvement and Practice, 4 (2):55–71, June 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Navarro:2005:SPM

[NvdH05] Emily Oh Navarro and André van der Hoek. Software process modeling for an educational software engineering simulation game. Software Process: Improvement and Practice, 10 (3):311–325, July 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Niazi:2006:CSF

[NWZ06] Mahmood Niazi, David Wilson, and Didar Zowghi. Critical success factors for software process improvement implementation: an empirical study. Software Process: Improvement and Practice, 11(2): 193–211, March 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ocampo:2005:SPC

[OBM05] Alexis Ocampo, Fabio Bella, and Jürgen Münch. Software process commonality analysis. Software Process: Improvement and Practice, 10(3):273– 285, July 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ojala:2008:IVA

[Oja08] Pasi Ojala. Implementing a value assessment for products: an industrial case study. Software Process: Improvement and Practice, 13(4):363–370, July 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ojala:2009:EVA

[Oja09] Pasi Ojala. Experiences of a value assessment for products. Software Process: Improvement and Practice, 14(1): 31–37, January 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ocampo:2009:RMS

[OM09] Alexis Ocampo and Jürgen Münch. Rationale modeling for software process evolution. Software Process: Improvement and Practice, 14(2): 85–105, March 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

OConchuir:2009:BGS

[ÓOÅF09] Eoin Ó Conchúir, Helena Holmström Olsson, Pär J. Ågerfalk, and Brian Fitzgerald. Benefits of global software development: exploring the unexplored. Software Process: Improvement and Practice, 14 (4):201–212, July 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic). See errata [Ano09].

Ovaska:2003:ACT

[ORM03] Päivi Ovaska, Matti Rossi, and Pentti Marttiin. Architecture as a coordination tool in multi-site software development. Software Process: Improvement and Practice, 8(4): 233–247, October 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Osterweil:2007:WWL

[Ost07] Leon J. Osterweil. What we learn from the study of ubiquitous processes. Software Process: Improvement and Practice, 12(5):399–414, September 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Setamanit:2007:USE

[oSWR07] Siri on Setamanit, Wayne Wakeland, and David Raffo.

Using simulation to evaluate global software development task allocation strategies. Software Process: Improvement and Practice, 12(5): 491–503, September 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ould:1996:CI

[Oul96] Martyn A. Ould. CMM and ISO 9001. Software Process: Improvement and Practice, 2 (4):281–289, December 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Padberg:2002:DSM

[Pad02] Frank Padberg. A discrete simulation model for assessing software project scheduling policies. Software Process: Improvement and Practice, 7(3–4):127–139, September/December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Padberg:2006:SOS

[Pad06] Frank Padberg. A study on optimal scheduling for software projects. Software Process: Improvement and Practice, 11 (1):77–91, January 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Prikladnicki:2003:GSD

[PAE03] Rafael Prikladnicki, Jorge Luis Nicolas Audy, and Roberto Evaristo. Global software development in practice lessons learned. Software Process: Improvement and Practice, 8(4):267–281, October 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Pfahl:2007:SDS

[PAER07] Dr Dietmar Pfahl, Ahmed Al-Emran, and Günther Ruhe. A system dynamics simulation model for analyzing the stability of software release plans. Software Process: Improvement and Practice, 12(5):475– 490, September 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Paulk:1996:EGE

[Pau96] Mark C. Paulk. Editorial:

Guest editorial for the special issue of papers from the 29th Hawaii International Conference on Systems Science. Software Process: Improvement and Practice, 2(3):165– 166, September 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Powell:1998:PSI

[PB98] Antony Powell and Duncan Brown. A practical strategy for industrial reuse improvement. Software Process: Improvement and Practice, 4(3): 173–182, September 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Paasivaara:2008:USG

[PDL08] Maria Paasivaara, Sandra Durasiewicz, and Casper Lassenius. Using scrum in a globally distributed project: a case study. Software Process: Improvement and Practice, 13 (6):527–544, November 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Penedo:2000:AWB

[Pen00] Maria H. Penedo. An active Web-based virtual room for small team collaboration. Software Process: Improvement and Practice, 5(4):251– 261, December 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

PhD:2009:EAM

[PhD09a] David M. Raffo PhD. Editorials: Addressing management issues. Software Process: Improvement and Practice, 14 (4):199–200, July 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

PhD:2009:EEP

[PhD09b] David M. Raffo PhD. Editorials: Examining process design and change. Software Process: Improvement and Practice, 14 (3):127–128, May 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Paasivaara:2003:CPG

[PL03] Maria Paasivaara and Casper Lassenius. Collaboration practices in global interorganizational software development projects. Software Process: Improvement and Practice, 8(4):183–199, October 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Poth:2007:SRE

[Pot07] Alexander Poth. SPI of the requirements engineering process for embedded systems using SPICE. Software Process: Improvement and Practice, 12 (6):579–584, November 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Poth:2009:PLR

[Pot09]  Alexander Poth.  Product

line requirements engineering in the context of process aspects in organizations with various domains. Software Process: Improvement and Practice, 14(6):315–323, November 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Pfahl:2002:IMI

[PR02] Dietmar Pfahl and Günther Ruhe. IMMoS: a methodology for integrated measurement, modelling and simulation. Software Process: Improvement and Practice, 7(3–4):189–210, September/ December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Pfahl:2005:GES

[PR05] Dietmar Pfahl and Ioana Rus. Guest editorials: Special issue on ProSim 2004, The 5th International Workshop on Software Process Simulation and Modeling, Edinburgh, Scotland, May 2004. Software Process: Improvement and Practice, 10(3):251–253, July 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1996:Ea

[PS96a] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 2(2):81, June 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1996:SCEa

[PS96b] Dewayne Perry and Wilhelm Schäfer. Short communications: Editorial. Software Process: Improvement and Practice, 2(1):1, March 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1996:SCEb

[PS96c] Dewayne Perry and Wilhelm Schäfer. Short communications: Editorial. Software Process: Improvement and Practice, 2(3):163, September 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1996:Eb

[PS96d] Dewayne Perry and Wilhelm Shäfer. Editorial. Software Process: Improvement and Practice, 2(4):237, December 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1997:Ea

[PS97a] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 3(1):1, March 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1997:Eb

[PS97b] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 3(2):69, June 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1997:Ec

[PS97c] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 3(3):139, September 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1997:Ed

[PS97d] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 3(4):199, December 1997. CODEN SPIPFL.

ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1998:Ea

[PS98a] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 4(1):1, March 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1998:Eb

[PS98b] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 4(2):53, June 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:1998:Ec

[PS98c] Dewayne Perry and Wilhelm Schäfer. Editorial. Software Process: Improvement and Practice, 4(3):99, September 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:2000:E

[PST00] Dewayne Perry, Wilhelm Schäfer, and Colin Tully. Editorial. Software Process: Improvement and Practice, 5(1): 1, March 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:2001:Ea

[PST01a] Dewayne Perry, Wilhelm Schäfer, and Colin Tully. Editorial. Software Process: Improvement and Practice, 6 (2):65, June 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:2001:Eb

[PST01b] Dewayne Perry, Wilhelm Schäfer, and Colin Tully. Editorial. Software Process: Improvement and Practice, 6(3): 123, September 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:2001:Ec

[PST01c] Dewayne Perry, Wilhelm Schäfer, and Colin Tully. Editorial. Software Process: Improvement and Practice, 6(4): 167, December 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:2002:Ea

[PST02a] Dewayne Perry, Wilhelm Schäfer, and Colin Tully. Editorial. Software Process: Improvement and Practice, 7(1): 1, March 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Perry:2002:Eb

[PST02b] Dewayne Perry, Wilhelm Schäfer, and Colin Tully. Editorial. Software Process: Improvement and Practice, 7 (2):45, June 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Porter:2006:TPI

[PYM^{+}06] Adam Porter, Cemal Yilmaz, Atif M. Memon, Arvind S. Krishna, Douglas C. Schmidt, and Aniruddha Gokhale. Techniques and processes for improving the quality and performance of open-source software. Software Process: Improvement and Practice, 11(2): 163–176, March 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Qin:2007:MPC

[Qin07] Shaowen Qin. Managing process change in software organizations: Experience and reflection. Software Process: Improvement and Practice, 12(5): 429–435, September 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Aagerfalk:2006:GES

[R06] Pär J. gerfalk and Jolita Ralyté. Guest editorials: Situational Requirements Engineering Processes: reflecting on method engineering and requirements practice. Software Process: Improvement and Practice, 11(5):447–450, September 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ramachandran:2005:CVA

[RA05] Muthu Ramachandran and Pat Allen. Commonality and variability analysis in industrial practice for product line improvement. Software Process:

Improvement and Practice, 10 (1):31–40, January 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ramachandran:2006:AIC

[Ram06] Muthu Ramachandran. Automated improvement for component reuse. Software Process: Improvement and Practice, 11(6):591–599, November 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Regev:2007:DBP

[RBW07] Gil Regev, Ilia Bider, and Alain Wegmann. Defining business process flexibility with the help of invariants. Software Process: Improvement and Practice, 12(1): 65–79, January 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Roehling:2000:SDM

[RCHSD00] Stephen T. Roehling, James S.

Collofello, Brian G. Hermann, and Dwight E. Smith-Daniels. System dynamics modeling applied to software outsourcing decision support. Software Process: Improvement and Practice, 5(2–3):169–182, June/September 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Raffo:2008:EIR

[RFoSS08] David Raffo, Robert Ferguson, Siri on Setamanit, and Bhuricha Sethanandha. Evaluating the impact of requirements analysis tools using simulation. Software Process: Improvement and Practice, 13(1): 63–73, January 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ramos:2008:SRP

[RGdASE08] Javier Saldaña Ramos, Javier García Guzmán, Antonio de Amescua Seco, and Ana Sanz Esteban. A study of reported practical experiences about TSP implementations. Software Process: Improvement and Practice, 13(5):397–409, September 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Rainer:2001:ASC

[RH01] Austen Rainer and Tracy Hall. An analysis of some ‘core studies’ of software process improvement. Software Process: Improvement and Practice, 6 (4):169–187, December 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Raffo:2000:CMM

[RHV00] David Raffo, Warren Harrison, and Joseph Vandeville. Coordinating models and metrics to manage software projects. Software Process: Improvement and Practice, 5(2–3):159–168, June/ September 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Richardson:2001:SPM

[Ric01] Ita Richardson. Software process matrix: a small company SPI model. Software Process: Improvement and Practice, 6 (3):157–165, September 2001. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Russwurm:2000:IEP

[RM00] Winfried Russwurm and Ludger Meyer. Integrated evaluation procedure for software/ hardware system development processes based on the Software Capability Maturity Model (CMM). Software Process: Improvement and Practice, 5(4):231–241, December 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Rejas-Muslera:2007:SLA

[RMCGSR07] Ricardo J. Rejas-Muslera, Juan J. Cuadrado-Gallego, Miguel-Angel Sicilia, and Daniel Rodríguez. SLA: a legal assurance process model for software engineering management. Software Process: Improvement and Practice, 12(2): 191–198, March 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Rolland:2009:MET

[Rol09] Colette Rolland. Method engineering: towards methods as services. Software Process: Improvement and Practice, 14 (3):143–164, May 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Rombach:1996:NNI

[Rom96] Prof. Dr. Dieter Rombach. News: New Institute for Applied Software Engineering Research. Software Process: Improvement and Practice, 2(2): 152–155, June 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Rout:2003:IIE

[Rou03] Terence P. Rout. ISO/IEC 15504 — evolution to an international standard. Software Process: Improvement and Practice, 8(1):27–40, January 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Rout:2007:GES

[Rou07]  Terence P. Rout.  Guest

editorials: Software process assessment and improvement: Selected papers from SPICE 2005. Software Process: Improvement and Practice, 12 (4):311–313, July 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Raffo:2008:GEG

[RPW08] David Raffo, Dietmar Pfahl, and Qing Wang. Guest editorials: Guest editors’ introduction. Software Process: Improvement and Practice, 13 (1):1–3, January 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ruiz:2004:IFS

[RRT04] Mercedes Ruiz, Isabel Ramos, and Miguel Toro. An integrated framework for simulationbased software process improvement. Software Process: Improvement and Practice, 9 (2):81–93, April 2004. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic). See correction [Ano05].

Raffo:2000:E

[RS00] D. Raffo and W. Scacchi.

Editorial. Software Process: Improvement and Practice, 5 (2–3):87–89, June/September 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Ramil:2002:QSM

[RS02] Juan F. Ramil and Neil Smith.

Qualitative simulation of models of software evolution. Software Process: Improvement and Practice, 7(3–4):95–112, September/December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Regev:2005:GEC

[RSB05] Gil Regev, Pnina Soffer, and Ilia Bider. Guest editorials: Creating and maintaining the fit between business processes and support systems. Software Process: Improvement and Practice, 10(4): 359–361, October 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Raatikainen:2005:CCS

[RSMM05] Mikko Raatikainen, Timo Soininen, Tomi Männistö, and Antti Mattila. Characterizing configurable software product families and their derivation. Software Process: Improvement and Practice, 10(1): 41–60, January 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Rout:2007:HII

[RT07] Terence P. Rout and Angela Tuffley. Harmonizing ISO/ IEC 15504 and CMMI. Software Process: Improvement and Practice, 12(4):361–371, July 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Riel:2009:IES

[RTM09] Andreas Riel, Serge Tichkiewitch, and Richard Messnarz. Integrated engineering skills for improving the system competence level. Software Process: Improvement and Practice, 14 (6):325–335, November 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Salo:2007:IIP

[SA07] Outi Salo and Pekka Abrahamsson. An iterative improvement process for agile software development. Software Process: Improvement and Practice, 12(1):81–100, January 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Saiedian:2003:PRM

[Sai03] Hossein Saiedian. Practical recommendations to minimize software capability evaluation risks. Software Process: Improvement and Practice, 8(3): 145–156, July 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Salvaneschi:2009:MKI

[Sal09] Paolo Salvaneschi. Managing knowledge for information system evolution: the MinimalEDoc methodology. Software Process: Improvement and Practice, 14(6):337–347, November 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Satriani:1997:NEN

[Sat97]  Giuseppe Satriani.  News:

ESI news: a European software best practice repository. Software Process: Improvement and Practice, 3(4):243– 245, December 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Schmid:2005:SMS

[SB05] Klaus Schmid and Stefan Biffl.

Systematic management of software product lines. Software Process: Improvement and Practice, 10(1):61–76, January 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Siakas:2006:SOQ

[SB06] Kerstin V. Siakas and Bo Balstrup. Software outsourcing quality achieved by global virtual collaboration. Software Process: Improvement and Practice, 11(3):319–328, May 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Schalken:2006:MDL

[SBvV06] Joost Schalken, Sjaak Brinkkemper, and Hans van Vliet. A method to draw lessons from project postmortem databases. Software Process: Improvement and Practice, 11(1):35– 46, January 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Scacchi:2000:USP

[Sca00] Walt Scacchi. Understanding software process redesign using modeling, analysis and simulation. Software Process: Improvement and Practice, 5(2–3):183–195, June/ September 2000. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Smith:2006:ABS

[SCFR06] Neil Smith, Andrea Capiluppi, and Juan Fernández-Ramil. Agent-based simulation of open source evolution. Software Process: Improvement and Practice, 11(4):423–434, July 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Schmidt:2007:SCO

[Sch07] Rainer Schmidt. Sercomp:

a component oriented method for flexible design and support of interorganizational service processes. Software Process: Improvement and Practice, 12(1):7–20, January 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Smith:2005:SOS

[SCR05] Neil Smith, Andrea Capiluppi, and Juan F. Ramil. A study of open source software evolution data using qualitative simulation. Software Process: Improvement and Practice, 10 (3):287–300, July 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Sora:2005:MVS

[ŞCVB05] Ioana Şora, Vladimir Creţu, Pierre Verbaeten, and Yolande Berbers. Managing variability of self-customizable systems through composable components. Software Process: Improvement and Practice, 10(1): 77–95, January 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Sechser:2009:MTP

[Sec09] Bernhard Sechser. The marriage of two process worlds. Software Process: Improvement and Practice, 14(6):349– 354, November 2009. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Simon:1997:RII

[SER^{+}97] Jean-Martin Simon, Khaled El Emam, Sonia Rousseau, Eric Jacquet, and Frederic Babey. The reliability of ISO/IEC PDTR 15504 assessments. Software Process: Improvement and Practice, 3(3): 177–188, September 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Suscheck:2008:JIL

[SF08] Charles A. Suscheck and Randal Ford. Jazz improvisation as a learning metaphor for the scrum software development methodology. Software Process: Improvement and Practice, 13(5):439–450, September 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Scacchi:2006:GEU

[SFF^{+}06] Walt Scacchi, Joseph Feller, Brian Fitzgerald, Scott Hissam, and Karim Lakhani. Guest editorials: Understanding free/ open source software development processes. Software Process: Improvement and Practice, 11(2):95–105, March 2006.

CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Stewart:2006:MRD

[SG06] Katherine J. Stewart and Sanjay Gosain. The moderating role of development stage in free/open source software project performance. Software Process: Improvement and Practice, 11(2):177–191, March 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Salvaneschi:2006:RPI

[SGB06] Paolo Salvaneschi, Daniele Grasso, and Maurizio Besurga. The role of process improvement in IT departments: experiences and lessons learned. Software Process: Improvement and Practice, 11(3):309– 317, May 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Singh:1996:ISI

[Sin96] Raghu Singh. International Standard ISO/IEC 12207 Software Life Cycle Processes. Software Process: Improvement and Practice, 2(1):35–50, March 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Stelzer:1998:SFO

[SM98] Dirk Stelzer and Werner Mellis. Success factors of organizational change in software process improvement.

Software Process: Improvement and Practice, 4(4):227– 250, December 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Scholtz:2003:UCD

[SM03] Jean Scholtz and Emile Morse. Using consumer demands to bridge the gap between software engineering and usability engineering. Software Process: Improvement and Practice, 8 (2):89–98, April 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Stelzer:1996:SPI

[SMH96] Dirk Stelzer, Werner Mellis, and Georg Herzwurm. Software process improvement via ISO 9000? Results of two surveys among European software houses. Software Process: Improvement and Practice, 2(3): 197–210, September 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Smite:2006:GSD

[Šmi06] Darja Šmite. Global software development projects in one of the biggest companies in Latvia: is geographical distribution a problem? Software Process: Improvement and Practice, 11(1):61–76, January 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Sneed:1998:MRL

[Sne98] Harry M. Sneed. Measuring reusability of legacy software systems. Software Process: Improvement and Practice, 4(1):43–48, March 1998. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Soffer:2005:SAI

[Sof05] Pnina Soffer. Scope analysis: identifying the impact of changes in business process models. Software Process: Improvement and Practice, 10(4): 393–402, October 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Spork:2008:EPD

[SP08] Gunther Spork and Uwe Pichler. Establishment of a performance driven improvement programme. Software Process: Improvement and Practice, 13 (4):371–382, July 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Scacchi:2002:ESI

[SR02] Walt Scacchi and David Raffo.

Editorials: Special issue on software process simulation and modeling. Software Process: Improvement and Practice, 7(3–4):91–93, September/December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Sanders:2007:RLT

[SR07] Marty Sanders and Ita Richardson. Research into longterm improvements in smallto medium-sized organisations using SPICE as a framework for standards. Software Process: Improvement and Practice, 12(4):351–359, July 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Steinmann:1996:STS

[SS96] Christian Steinmann and Hans Stienen. SynQuest — tool support for software selfassessments. Software Process: Improvement and Practice, 2 (1):5–12, March 1996. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Schneider:2001:MEB

[SS01] Kurt Schneider and Thilo Schwinn. Maturing experience base concepts at DaimlerChrysler. Software Process: Improvement and Practice, 6 (2):85–96, June 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Siakas:2007:APC

[SS07] Kerstin V. Siakas and Errikos Siakas. The agile professional culture: a source of agile quality. Software Process: Improvement and Practice, 12(6): 597–610, November 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Sawyer:1997:RPI

[SSV97] Pete Sawyer, Ian Sommerville, and Stephen Viller. Requirements process improvement through the phased introduction of good practice. Software Process: Improvement and Practice, 3(1):19–34, March 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Savioja:2007:MPF

[ST07] Erkki Savioja and Markku Tukiainen. Measurement practices in financial software industry. Software Process: Improvement and Practice, 12(6): 585–595, November 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Stallinger:2000:SPS

[Sta00] Friedrich Stallinger. Software process simulation to support ISO/IEC 15504 based software process improvement. Software Process: Improvement and Practice, 5(2–3):197–209, June/September 2000. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Snowdon:2007:AFF

[SWG^{+}07] R. A. Snowdon, B. C. Warboys, R. M. Greenwood, C. P. Holland, P. J. Kawalek, and D. R. Shaw. On the architecture and form of flexible process support. Software Process: Improvement and Practice, 12 (1):21–34, January 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Taxen:2005:SAT

[Tax05] Lars Taxén. A sociotechnical approach towards alignment. Software Process: Improvement and Practice, 10(4): 427–439, October 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Tomaszewski:2008:RPT

[TBD08] Piotr Tomaszewski, Patrik Berander, and Lars-Ola Damm. Regular papers: From traditional to streamline development — opportunities and challenges. Software Process: Improvement and Practice, 13 (2):195–212, March 2008. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Turetken:2008:PMP

[TD08] Oktay Turetken and Onur Demirors. Process modeling by process owners: a decentralized approach. Software Process: Improvement and Practice, 13(1):75–87, January 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Taylor:2008:PSS

[TGC^{+}08] Philip S. Taylor, Des Greer, Gerry Coleman, Kevin Mc-Daid, and Frank Keenan. Preparing small software companies for tailored agile method adoption: Minimally intrusive risk assessment. Software Process: Improvement and Practice, 13(5):421–437, September 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Tripathi:2008:SOS

[TGG08] A. K. Tripathi, Ratneshwer Gupta, and Manjari Gupta. Some observations on software processes for CBSE. Software Process: Improvement and Practice, 13(5):411–419, September 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Tuffley:2004:SSO

[TGM04] Angela Tuffley, Bill Grove, and Gary McNair. SPICE for small organisations. Software Process: Improvement and Practice, 9(1):23–31, January 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Thorwart:1998:API

[Tho98] K. Thorwart. The AMETIST process improvement experiment: towards efficient team development in small companies. Software Process: Improvement and Practice, 4(1): 11–18, March 1998. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Thomson:1997:ASP

[TM97] Helen E. Thomson and Pam Mayhew. Approaches to software process improvement.

Software Process: Improvement and Practice, 3(1):3–17, March 1997. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Tasse:2001:VBP

[TM01] Josée Tassé and Nazim H. Madhavji. View-based process elicitation: a user’s perspective. Software Process: Improvement and Practice, 6(3): 125–139, September 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Thu:2005:TPC

[TNT^{+}05] Tran Dan Thu, Tran Hanh Nhi, Dong Thi Bich Thuy, Bernard Coulette, and Xavier Cregut. Topological properties for characterizing wellformedness of process components. Software Process: Improvement and Practice, 10(2): 217–247, April 2005. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Tervonen:1996:RIP

[TOK96] Ilkka Tervonen and Harri Oinas-Kukkonen. Reorganizing the inspection process: problems encountered and resolved. Software Process: Improvement and Practice, 2(2): 97–110, June 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Tomes:1996:IQS

[Tom96]  Dr Nils Tomes.  Interna-

tional quality: Software process improvement approaches in Japan and the UK. Software Process: Improvement and Practice, 2(2):83–95, June 1996. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Tsumaki:2006:FMR

[TT06] Toshihiko Tsumaki and Tetsuo Tamai. Framework for matching requirements elicitation techniques to project characteristics. Software Process: Improvement and Practice, 11 (5):505–519, September 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Vakaslahti:1997:PIF

[Vak97] Pasi Vakaslahti. Process improvement frameworks — a small case study with People Capability Maturity Model. Software Process: Improvement and Practice, 3(4):225– 234, December 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

vandeWeerd:2006:SIM

[vdWBSV06] Inge van de Weerd, Sjaak Brinkkemper, Jurriaan Souer, and Johan Versendaal. A situational implementation method for Web-based content management system-applications: method engineering and validation in practice. Software Process: Improvement and Practice, 11(5):521–538, September 2006. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Verlage:1997:NES

[Ver97] Martin Verlage. News: Experience with software process modeling. Software Process: Improvement and Practice, 3 (2):133–136, June 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

vanOmmering:2005:ESI

[vOW05] Rob van Ommering and David Weiss. Editorials: Special issue on SPLC 2004, The 3rd Software Product Line Conference, Boston, MA, August– September 2004. Software Process: Improvement and Practice, 10(2):101–102, April 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

vonWangenheim:2006:SBS

[vWVS06] Christiane Gresse von Wangenheim, Timo Varkoi, and Clênio F. Salviano. Standard based software process assessments in small companies. Software Process: Improvement and Practice, 11(3): 329–335, May 2006. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Walker:2003:LPC

[Wal03] A. J. Walker. Level 5 process capability achievement: a case study from Software Engineering Research Management. Software Process: Improvement and Practice, 8(1): 51–62, January 2003. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wilkie:2007:LOM

[WCM^{+}07] F. G. Wilkie, F. Mc Caffery, D. McFall, Neil Lester, and Emmanuel Wilkinson. A Low-overhead method for software process appraisal. Software Process: Improvement and Practice, 12(4):339–349, July 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wernick:2007:ESS

[Wer07]  Paul Wernick.  Editorials:

SPIP special issue on SPW/ ProSim 2006. Software Process: Improvement and Practice, 12(5):373–375, September 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Weske:2001:AMI

[WGHS01] Mathias Weske, Thomas Goesmann, Roland Holten, and Rüdiger Striemer. Analysing, modelling and improving workflow application development processes. Software Process: Improvement and Practice, 6 (1):35–46, March 2001. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wang:2008:EFE

[WGJ^{+}08] Qing Wang, Lang Gou, Nan Jiang, Meiru Che, Ronghui Zhang, Yun Yang, and Mingshu Li. Estimating fixing effort and schedule based on defect injection distribution. Software Process: Improvement and Practice, 13(1):35–50, January 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wernick:2002:SGS

[WH02] Paul Wernick and Tracy Hall. Simulating global software evolution processes by combining simple models: an initial study. Software Process: Improvement and Practice, 7(3–4):113–126, September/December 2002. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wieczorek:1997:NES

[Wie97] Isabella Wieczorek. News: On the establishment of successful measurement programs in industry. Software Process: Improvement and Practice, 3(3): 191–194, September 1997. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wilkie:2005:ECP

[WMM05] F. G. Wilkie, D. McFall, and F. McCaffery. An evaluation of CMMI process areas for smallto medium-sized software development organisations. Software Process: Improvement and Practice, 10(2):189–201, April 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wakeland:2004:UDE

[WMR04] Wayne W. Wakeland, Robert H.

Martin, and David Raffo. Using design of experiments, sensitivity analysis, and hybrid simulation to evaluate changes to a software development process: a case study. Software Process: Improvement and Practice, 9(2):107–119, April 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wolf:2008:DDS

[WND08] Timo Wolf, Thanh Nguyen, and Daniela Damian. Does distance still matter? Software Process: Improvement and Practice, 13(6):493–510, November 2008. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wernick:2004:ESI

[WS04] Paul Wernick and Walt Scacchi. Editorials: Special issue on ProSim 2003, The 4th International Workshop on Software Process Simulation and Modeling, Portland, OR, May 2003. Software Process: Improvement and Practice, 9(2):51–53, April 2004. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Wakeland:2005:HOV

[WSR05] Wayne Wakeland, Stephen Shervais, and David Raffo. Heuristic optimization as a V&V tool for software process simulation models. Software Process: Improvement and Practice, 10(3):301–309, July 2005. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Xiao:2007:ALJ

[XOZ^{+}07] Junchao Xiao, Leon J. Osterweil, Lei Zhang, Alexander Wise, and Qing Wang. Applying Little-JIL to describe process–agent knowledge and support project planning in SoftPM. Software Process: Improvement and Practice, 12(5): 437–448, September 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Yang:2007:IPD

[YB07] Ye Yang and Barry Boehm. Improving process decisions in COTS-based development via risk-based prioritization. Software Process: Improvement and Practice, 12(5):449–460, September 2007. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Yilmaz:2007:ITE

[YP07] Levent Yilmaz and Jared Phillips. The impact of turbulence on the effectiveness and efficiency of software development teams in small organizations. Software Process: Improvement and Practice, 12 (3):247–265, May 2007. CO-DEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Zopf:2009:SFG

[Zop09] Siegfried Zopf. Success factors for globally distributed projects. Software Process: Improvement and Practice, 14 (6):355–359, November 2009. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).